### v2

# Permutation twisted cohomology, remixed

Abstract.
We generalize Balmer and Gallauer’s (permutation) twisted cohomology ring, working toward an alternative method of deducing the Balmer spectrum of the derived category of permutation modules for any finite $p$-group. The construction comes equipped with a canonical comparison map from the Balmer spectrum to the homogeneous spectrum of the twisted cohomology ring. We show the comparison map is injective for any finite $p$-group and furthermore, an open immersion when the twisted cohomology ring is noetherian. For elementary abelian $p$-groups, the twisted cohomology ring coincides with Balmer and Gallauer’s original construction.
To perform this construction, we utilize endotrivial complexes (i.e. the invertible objects of the derived category of permutation modules) arising up to a shift from Bredon homology of representation spheres. This topological structure allows us to construct certain $p$-local isomorphisms, from which we build a refined open cover of the Balmer spectrum indexed by conjugacy classes of subgroups of $G$. Under this open cover, every endotrivial is isomorphic to a shift of the tensor unit in each localization, thus verifying that all endotrivials are line bundles. When the twisted cohomology ring is noetherian, this open cover endows the Balmer spectrum with Dirac scheme structure.
Key words and phrases:
permutation module, twisted cohomology, comparison map, Balmer spectrum, tensor-triangular geometry
2020 Mathematics Subject Classification:
18G80, 18G90, 18M05, 20C20, 20J05
Introduction
Let $k$ be a field of prime characteristic $p$ and $G$ be a finite group. Permutation $kG$-modules (modules which admit a $G$-stable $k$-basis), and their direct summands, $p$-permutation kGkG-modules, are perhaps the second easiest class of representations of $G$ over $k$ to define. In a wildly vast representation-theoretic cosmos of $kG$-modules, our $p$-permutation modules may seem tiny and insignificant, drowning in a sea of symmetries. However, our protagonists, or more precisely, the bounded homotopy category of $kG$-modules $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm}(kG))$, exhibits significant control over the representation theory of $kG$ in general. For instance, if $G$ has an abelian Sylow $p$-subgroup $S$, Broué’s abelian defect group conjecture predicts a triangulated equivalence $\operatorname{K}_{b}(p\operatorname{-perm}(kGb_{0}))\cong\operatorname{K}_{b}(p\operatorname{-perm}(kN_{G}(S)c_{0}))$ where $b_{0}$ and $c_{0}$ denote the principal blocks of $kG$ and $kN_{G}(S)$ respectively. Perhaps more surprisingly, Balmer and Gallauer recently demonstrated that $\mathscr{K}(G)$ admits the bounded derived category $\operatorname{D}_{b}(\text{mod}(kG))$ as a Verdier quotient [BG23a], strengthening an unpublished result of Rouquier [Rou06]. Although these $p$-permutation are few, they are mighty; together, they recover the entire representation theory of $kG$.
If the reader is less representation-theoretically inclined, fear not! Our $p$-permutation category $\mathscr{K}(G)$ is the compact part of a “big” (i.e. rigidly-compactly generated) tt-category $\mathscr{D}(G):=\operatorname{DPerm}(G;k)$, the derived category of permutation modules. This category $\mathscr{D}(G)$ is equivalent to numerous big tt-categories arising from seemingly disparate areas of mathematics, such as:
  1. (a)
    the derived category of cohomological $k$-linear Mackey functors over $G$;
  2. (b)
    the homotopy category of modules over the constant Green functor $\operatorname{H}\underline{k}$ in genuine $G$-spectra;
  3. (c)
    the category of $k$-linear Artin motives generated by motives of intermediate fields in a Galois extension with Galois group $G$.
See [BG23b] for an overview of these correspondences; we remark Fuhrmann recently ascended equivalence (b) to the $\infty$-categorical level [Fuh25]. Hence any results concerning $\mathscr{D}(G)$ and $\mathscr{K}(G)$ translate over to the aforementioned categories and their compact parts respectively.
Continuing their foray into the land of permutation modules, Balmer-Gallauer deduced the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ of $\mathscr{K}(G)$ in [BG25], a landmark achievement. Their construction of permutation twisted cohomology, a modification of the usual cohomology ring, plays a key role in this classification. The authors consider morphisms from the tensor unit to shifts of certain invertible objects (i.e. endotrivial complexes) corresponding to subgroups $N\triangleleft G$ of index $p$. The collection of these morphisms forms a multigraded ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, the twisted cohomology ring, extending the usual notion of a cohomology ring of a triangulated category, and comes with a canonical comparison map,
$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)),$
which was previously constructed in [Bal10a]. A twist of the cohomology ring is necessary - since we are considering a homotopy category, the usual cohomology ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)}(k)$ is a tad drab on its own. The miracle of this construction is:
Theorem.
[BG25, Theorem 10.5] Let $E$ be a finite elementary abelian $p$-group. The comparison map $\operatorname{comp}_{E}\colon\operatorname{Spc}(\mathscr{K}(E))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E))$ identifies $\operatorname{Spc}(\mathscr{K}(E))$ with an open subspace of the homogeneous spectrum of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$.
This remarkable fact, along with a colimit theorem [BG25, Theorem 11.10] reducing to the case of $G$ elementary abelian á la Quillen stratification, completes the description of $\operatorname{Spc}(\mathscr{K}(G))$ as a topological space, and as a Dirac scheme in the sense of [HP23]. As a set, $\operatorname{Spc}(\mathscr{K}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}\operatorname{Spc}(\operatorname{D}_{b}(kG))$ ([BG25, Theorem 7.16]); we set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$. Hausmann and Schwede further considered the twisted cohomology ring for elementary abelian $2$-groups [HS25] (in the Mackey functor setting), which they entitled “Representation-graded Bredon homology.” The authors determined a minimal generating set and proved that for $p=2$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(E)$ is nilpotent-free.
Separately, the author deduced the Picard group of $\mathscr{K}(G)$, i.e. the group of endotrivials, in a sequence of papers [Mil24, Mil25b, Mil25a]. Significant parts of the classification were also completed by Bachmann in his dissertation [Bac16] in the context of Artin motives. When $G$ is a finite $p$-group, morally the endotrivials arise from Bredon homology of representation spheres, and therefore have additional topological structure. This was also observed by Yalçin in the context of $G$-Moore spaces [Yal17], where an analogous classification result was determined. Therefore, one obtains a canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ associated to real representations modulo Adams operations. Interestingly, for elementary abelian $p$-groups, this basis corresponds exactly to the endotrivials used by Balmer-Gallauer in their twisted cohomology ring. The natural question for one to ask is:
Question.
Can we recover the Balmer spectrum $\operatorname{Spc}(\mathscr{K}(G))$ by further twisting (or if you will, re-mixing) $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ by the endotrivial complexes arising (up to shifts) from representation spheres?
We consider twisting by shifts of complexes coming from genuine $\mathbb{R}G$-modules as opposed to virtual representations, as if we were to twist by all endotrivials, the graded ring becomes unmanageably large (c.f. [BG25, Remark 12.24]). The analogous comparison to make here is that the group cohomology ring $\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ is noetherian and recovers the Balmer spectrum of the derived category $\operatorname{D}_{b}(kG)$, whereas the Tate cohomology ring is too large. We call such endotrivials effective, and their corresponding h-mark homomorphisms are monotonically decreasing superclass functions. Similarly, Hausmann-Schwede call the corresponding “quadrant” of twisted cohomology associated to twists by effective endotrivials the effective cone.
Denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$ arising from nontrivial irreducible real representations of $G$ by $\mathcal{B}(G)$. Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e. tuples of non-negative integers indexed by the set $\mathcal{B}(G)$. The monoid of twists is equivalently nothing more than the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$. Then the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
$\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)=\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G;k):=\bigoplus_{s\in\mathbb{Z}}\bigoplus_{q\in\mathbb{N}^{\mathcal{B}(G)}}\operatorname{Hom}_{\mathscr{K}(G)}\left(k,\bigotimes_{C\in\mathcal{B}(G)}C^{\otimes q(C)}[s]\right)$
is the (re-)twisted cohomology ring; see Definition 5.1. We note that only non-positive shifts $s\leq 0$ play a nontrivial role.
We take the first steps towards answering this question for finite $p$-groups by proving injectivity of the comparison map. The $p$-group case is perhaps the most fundamental: given a finite group $G$ with $p$-Sylow $S$, one can determine $\operatorname{Spc}(\mathscr{K}(G))$ from $\operatorname{Spc}(\mathscr{K}(S))$ via restriction.
Theorem A. (Theorem 6.1).
Let $G$ be a finite $p$-group. The comparison map $\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is injective.
Generally, injectivity of the comparison map can be a harder condition to prove (if it holds at all, which it frequently may not!) than surjectivity (after a categorical localization), which general tt-geometry usually resolves. Circling back to Balmer-Gallauer’s original construction, another feature of $\operatorname{Spc}(\mathscr{K}(E))$ for $E$ elementary abelian is a spectral open cover $\{U(H)\}_{H\leq E}$ of $\operatorname{Spc}(\mathscr{K}(E))$ indexed by subgroups of $E$. This cover has two remarkable features:
  1. (a)
    Over each open $U(H)$, all endotrivials for $E$ are trivial, i.e. isomorphic to a shift of the tensor unit in the localization $\mathscr{L}(H):=\mathscr{K}(E)|_{U(H)}$;
  2. (b)
    There is a homeomorphism between each $U(H)$ and the homogeneous spectrum of the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}_{\mathscr{L}(H)}^{\scriptscriptstyle\bullet}(k)$ in the localization $\mathscr{L}(H)$. In particular, $\operatorname{Spc}(\mathscr{K}(E))$ is a Dirac scheme.
In addition, under this open cover, there is an identification $\mathscr{K}(E)|_{U(1)}\cong\operatorname{D}_{b}(kE)$ - in other words, $U(1)$ is nothing more than the “cohomological open,” corresponding to the image under $\operatorname{Spc}$ of the localization $\mathscr{K}(E)\twoheadrightarrow\operatorname{D}_{b}(kE)$. We construct a refined open cover with these same properties.
Theorem B.
Let $G$ be a finite $p$-group. Then there exists an open cover
$\{U(H)\}_{H\in s_{p}(G)/G}$ (Construction 4.5) indexed by conjugacy classes of subgroups for which the following holds.
  1. (a)
    Each open $U(H)$ contains a unique closed point of $\operatorname{Spc}(\mathscr{K})$, $\mathfrak{m}_{H}$, corresponding to the unique closed point of $V_{{G}/\!\!/{H}}$; (Corollary 4.11)
  2. (b)
    The open $U(1)$ is equivalently the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG))$ of $\operatorname{Spc}(\mathscr{K}(G))$. In other words, we have an equivalence of categories $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$; (Theorem 4.17)
  3. (c)
    Every endotrivial $C$ is a line bundle, i.e. over each open $U(H)$ is isomorphic to a shift of the tensor unit in the localization $\mathscr{K}(G)|_{U(H)}$. In particular, we have an isomorphism $C\cong k[h_{C}(H)]$ where $h_{C}$ denotes the unique h-mark function associated to $C$. (Theorem 4.13)
Finally, if $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, we completely recover the Balmer spectrum.
Theorem C. (Corollaries 6.4, 6.5).
If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, the comparison map is an open immersion and for each subgroup $H\leq G$, restricts to a local homeomorphism
$\operatorname{comp}_{\mathcal{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\operatorname{End}_{\mathscr{L}_{G}(H)}^{\scriptscriptstyle\bullet}(k)).$
In this case $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme.
To obtain these results, it is seemingly necessary to twist by all effective endotrivials (or at minimum, a full-rank subset); see for instance Remark 4.12. In this sense, we believe our construction for $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is the “correct” generalization, and not too big. The not-so-secret sauce to the above results is a generalization of Balmer-Gallauer’s maps $a_{N}$ and $b_{N}$ for any effective endotrivial. For any effective endotrivial $C$, we construct unique (up to scaling) maps $\iota_{C}^{H}$ for every subgroup $H\leq G$ for which $\Psi^{H}(\iota_{C}^{H})$ is a quasi-isomorphism. Proving these maps exist is rather technical and relies heavily on the topological origins of effective endotrivials. These maps may not exist for chain complexes not arising from chain complexes of free modules over the orbit category $\Gamma_{G}$ (in particular, non-effective endotrivials), see Remark 3.8.
Theorem D. (Theorem 4.2).
Let $G$ be a finite $p$-group. For every effective endotrivial complex (see Definition 2.1) and subgroup $H\leq G$, there exists a homomorphism $\iota^{H}_{C}:k\to C[h_{C}(H)]$ such that $\Psi^{H}(\iota^{H}_{C})$ is a quasi-isomorphism.
Note that there are numerous maps that satisfy the above property (see Remark 4.3); our maps $\iota^{H}_{C}$ satisfy additional technical conditions. Finally, we prove some technical results regarding these chain complexes. Although they are not strictly necessary for our constructions, they contribute to our understanding of effective endotrivials. These propositions may be known to experts, but we reprove them for completeness.
Theorem E. (Proposition 3.11, Corollary 3.12).
Let $G$ be a finite group. Suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. Then $C$ has a contractible direct summand if and only if $C^{?}$ does. In particular, if $C$ is an indecomposable effective endotrivial, then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
Open questions
There is but one essential ingredient missing to this story, that of noetherianity of the remixed twisted cohomology ring $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. At present, we conjecture that $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, and in fact, generated by homomorphisms $k\to C[s]$ for $C$ an irreducible endotrivial. Sadly, the strategy of [BG25, Lemma 12.12] to prove noetherianity seems to not carry over to the general setting. We (sketchily) remark Balmer-Gallauer’s proof can be adapted to any finite abelian $p$-group, but we omit a proof, as it follows near identically as in the elementary abelian case.
A new possibility we propose is comparing twisted cohomology to the usual cohomology ring. We construct a homomorphism from the twisted cohomology ring to a direct product of $p$-local group cohomology rings (see Definition 5.5), and whose kernel is nilpotent. Any group cohomology ring is well-known to be noetherian [Ven59, Eve61], therefore if the image of $\prod\hat{\Psi}$ is finite in each $p$-local group cohomology ring, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian modulo nilpotents. See Remark 6.6 for details.
Beyond this missing link, there remain numerous questions left to answer. First, it is quite reasonable to ask if a generalization of twisted cohomology exists for all finite groups. Not all endotrivials $C$ arise from a representation spheres (or any reasonable topological space), or can be expressed as (up to homotopy) a chain complex of permutation modules (see e.g. [BG23a, Corollary 5.6]). As a consequence, many of our maps $\iota^{H}_{C}$ cannot be constructed outside of the $p$-group setting; circumventing this issue is the main challenge. To go about constructing twisted cohomology, one may first have to determine $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$, which at present is unknown (for non-$p$-groups). We ask if $\operatorname{Pic}(\operatorname{K}_{b}(\operatorname{perm}(kG)))$ and the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ generated by representation spheres are in fact equal, and if they are finite index subgroups of $\operatorname{Pic}(\mathscr{K}(G))$.
Second, while our twisted cohomology ring may be sufficient to capture the subtleties of $\operatorname{Spc}(\mathscr{K}(G))$, it is obviously more cumbersome. An explicit presentation in terms of generators and relations is at present seemingly beyond our reach. Even explicitly writing down endotrivials for larger groups (e.g. $p$-rank at least 3) is a significant challenge and relatively unexplored. Finally, although every endotrivial is a line bundle, it is currently unknown and an interesting question whether $\operatorname{Pic}(\mathscr{K}(G)|_{U(H)})\cong\mathbb{Z}$; we expect this to be the case. In correspondence, Gallauer gave a proof sketch of this for $p=2$ and $G$ elementary abelian, but not much is known beyond this case.
Organization
The paper is organized as follows. Section 1 covers some quick preliminaries regarding $p$-permutation modules, modular fixed points/Brauer quotients, and endotrivial complexes. Section 2 further reviews Borel-Smith functions. Section 3 covers chain complexes of free modules over the orbit category and effective endotrivials, and proves some technical lemmas. Section 4 constructs the local quasi-isomorphisms $\iota^{H}_{C}$ and the open cover $\{U(H)\}$ of the Balmer spectrum. Section 5 constructs the twisted cohomology ring. Section 6 proves injectivity of the comparison map and corollaries, generalizing the results of Balmer-Gallauer in the noetherian setting, and discusses noetherianity.
Notation
Our notation mostly follows [BG25], as opposed to [Mil25a]. Any group $G$ is assumed to be finite. The Weyl group G//H{G}/\!\!/{H} of $H\leq G$ is the subquotient $N_{G}(H)/H$. Given two subgroups $H,K\leq G$, we write $H=_{G}K$ if $H,K$ are $G$-conjugate, and $H\leq_{G}K$ to denote $H$ is a subgroup of a $G$-conjugate of $K$. We denote the set of all $p$-subgroups of $G$ by $s_{p}(G)$, and write $s_{p}(G)/G$ to denote a choice of conjugacy class representatives. We write $\operatorname{CF}(G)$ to denote the additive group of superclass functions, and if $G$ is not a $p$-group, we write $\operatorname{CF}(G,p)$ to denote the group of superclass functions valued on $p$-subgroups. We denote the trivial $kG$-module, the tensor unit of $\mathscr{K}(G)$, as $k$. Given a module $M$, $M[i]$ denotes the chain complex with $M$ in homological degree $i\in\mathbb{Z}$ and zero in all other degrees. [1]-[1] also denotes the shift functor on $\mathscr{K}(G)$. We denote by $\operatorname{H}^{\scriptscriptstyle\bullet}(G)=\operatorname{H}^{\scriptscriptstyle\bullet}(G;k)$ the $\mathbb{N}$-graded cohomology ring of the tensor unit $k$ in $\operatorname{D}_{b}(kG):=\operatorname{D}_{b}(kG\operatorname{-mod})$ (with both odd and even-degree shifts for $p$ odd).
Acknowledgments
I am indebted to Paul Balmer and Martin Gallauer for numerous conversations and unwieldy email threads about their work and tensor-triangular geometry in abstracto. I also am grateful to Robert Boltje for his mentorship and suggestions, those of which heavily influenced Section 3 in particular, and the UC Santa Cruz Department of Mathematics for their generous graduate student travel support, which enabled conversations leading to the ideas in this paper to occur.
1. Preliminaries
We expediently review some of the essentials of permutation modules from [BG25] and endotrivial complexes [Mil25a]. For an overview of tensor-triangular geometry, we refer the reader to [Bal05, Bal10b], and for an in-depth “classical” overview of $p$-permutation modules and the Brauer quotient, we refer the reader to [Lin18, Chapter 5] or [Las23].
Definition 1.1.
A $kG$-module $M$ is:
  1. (a)
    a permutation module if $M$ admits a $G$-stable basis, or equivalently, $M\cong kX$ for some $G$-set $X$;
  2. (b)
    a $p$-permutation module (where $p$ denotes the characteristic of $k$) if $M$ is a direct summand of a permutation module, or equivalently, $\operatorname{Res}^{G}_{S}M$ is a permutation module for some Sylow $p$-subgroup $S$ of $G$. In particular, if $G$ is a finite $p$-group, every $p$-permutation module is a permutation module.
We set $\mathscr{K}(G):=\operatorname{K}_{b}(p\operatorname{-perm})$, the bounded homotopy category of $p$-permutation modules.
Definition 1.2.
Let $K\leq H\leq G$ be subgroups of $G$. The augmentation homomorphism for $G$-sets $\operatorname{aug}\colon G/K\to G/H$ is the homomorphism induced by the assignment $1K\mapsto 1H$. The augmentation homomorphism between the permutation modules $\operatorname{aug}\colon k[G/K]\to k[G/H]$ is the induced map on vector spaces. We note the standard example of the augmentation homomorphism is when $H=G$ and $K=1$.
We have an analogous coaugmentation homomorphism $\operatorname{coaug}\colon k[G/H]\to k[G/K]$ induced from the assignment $1H\mapsto\sum_{g\in[H/K]gK}$. This homomorphism of permutation modules is in no way induced from any $G$-set homomorphism.
Remark 1.3.
By [BG23a], we have a surjective tt-functor $\mathscr{K}(G)\twoheadrightarrow\operatorname{D}_{b}(kG)$ with kernel $\mathscr{K}_{ac}(G)$, the tt-ideal consisting of all acyclic complexes. Set $V_{G}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, then [BG25, Proposition 3.22] asserts that this functor induces an open inclusion $V_{G}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. We say $V_{G}\subseteq\operatorname{Spc}(\mathscr{K}(G))$ is the cohomological open. As noted in the introduction of [BG25], the crux of the matter is its closed complement, $\operatorname{supp}(\mathscr{K}_{ac})$.
The Brauer quotient is an explicit construction on the level of modules which mimics taking fixed points of $G$-sets. Balmer-Gallauer adapt this construction to work on the level of big categories, dubbing them modular fixed points functors.
Theorem 1.4.
[BG25, Proposition 2.7] For every $p$-subgroup $H\leq G$ there exists a coproduct-preserving tt-functor on the big derived category of permutation modules
$\Psi^{H}\colon\mathscr{D}(G)\to\mathscr{D}({G}/\!\!/{H})$
such that $\Psi^{H}(kX)\cong k[X^{H}]$ for every $G$-set $X$. In particular, this functor preserves compacts and restricts to a tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}({G}/\!\!/{H})$.
Notation 1.5.
Each tt-functor $\Psi^{H}$ induces a continuous map on spectra
$\operatorname{Spc}(\Psi^{H}):=\psi^{H}\colon\operatorname{Spc}(\mathscr{K}({G}/\!\!/{H}))\to\operatorname{Spc}(\mathscr{K}(G)),$
and composing with the surjection $\hat{\Psi}^{H}\colon\mathscr{K}(G)\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])$ induces another continuous map on spectra
$\hat{\psi}^{H}\colon V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G)).$
In fact, the Balmer spectrum, as a set, is built entirely from these “$p$-local” cohomological opens.
Theorem 1.6.
[BG25, Theorem 2.10] Every point of $\operatorname{Spc}(\mathscr{K}(G))$ is the image $\hat{\psi}^{H}(\mathfrak{p})$ of a point $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ for some subgroup $H\leq G$ unique up to $G$-conjugation, i.e. we have $\hat{\psi}^{H}(\mathfrak{p})=\hat{\psi}^{H^{\prime}}(\mathfrak{p}^{\prime})$ if and only if there exists a $g\in G$ such that $^{g}H=H^{\prime}$ and $^{g}\mathfrak{p}=\mathfrak{p}^{\prime}$.
Notation 1.7.
We denote by $\mathscr{P}_{G}(H,\mathfrak{p})\in\operatorname{Spc}(\mathscr{K})$ the image of the prime $\mathfrak{p}\in V_{{G}/\!\!/{H}}$ under $\hat{\psi}^{H}:V_{{G}/\!\!/{H}}\to\operatorname{Spc}(\mathscr{K}(G))$. By [BG25, Theorem 2.10], every prime is, up to $G$-conjugation, uniquely expressible in this way.
Theorem 1.8.
[BG25, Theorem 2.11] The family of functors
$\{\mathscr{D}(G)\xrightarrow{\Psi^{H}}\mathscr{D}({G}/\!\!/{H})\twoheadrightarrow\operatorname{K}(\operatorname{Inj}(k[{G}/\!\!/{H}]))\}_{H\in s_{p}(G)/G}$
indexed by conjugacy classes of $p$-subgroups $H\leq G$ is conservative. This restricts to a conservative family of functors $\{\mathscr{K}(G)\xrightarrow{\Psi^{H}}\mathscr{K}({G}/\!\!/{H})\twoheadrightarrow\operatorname{D}_{b}(k[{G}/\!\!/{H}])\}_{H\in s_{p}(G)/G}$ on compacts.
Next, we briefly review the classification of endotrivial complexes.
Definition 1.9.
A chain complex $C\in\mathscr{K}(G)$ is endotrivial if it is an invertible object, i.e. $C^{*}\otimes_{k}C\simeq k[0]$.
Remark 1.10.
By the Künneth formula for $kG$-modules, for any $p$-subgroup $H\leq G$, $\Psi^{H}(C)$ has nonzero homology in exactly one homological degree, with that homology having $k$-dimension one. Let $h_{C}(H)\in\mathbb{Z}$ denote the unique integer for which $\operatorname{H}_{h_{C}(H)}(\Psi^{H}(C))\neq 0$.
Proposition 1.11.
[Mil24, Definition 3.5]
  1. (a)
    The function $h_{C}\colon s_{p}(G)\to\mathbb{Z}$ is a well-defined superclass function, i.e. a function constant on conjugacy classes of subgroups of $G$.
  2. (b)
    The assignment $h\colon\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}(G,p)$ given by $C\mapsto h_{C}$ is a well-defined group homomorphism.
Given an endotrivial $C$, we call $h_{C}$ the h-mark function of $C$, and call the homomorphism $h$ the h-mark homomorphism. Under the hood, $h$ is nothing more than a numerical avatar of the conservative family of functors $\{\hat{\Psi}^{H}\}_{H\in s_{p}(G)}$. Since the invertible objects of $\operatorname{D}_{b}(kG)$ consists of shifts of $k$-dimension one $kG$-modules, $h_{C}$ simply tracks the image of $C$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$ where $H$ runs through all $p$-subgroups of $G$.
Definition 1.12.
A Borel-Smith function is a superclass function $f\in\operatorname{CF}(G,p)$ satisfying the following three conditions, which we call the Borel-Smith conditions.
  1. (a)
    If $p$ is odd, then for any subquotient $T/S$ of $G$ of order $p$, $f(T)\equiv f(S)\mod 2$.
  2. (b)
    If $p=2$, then for any sequence of subgroups $H\trianglelefteq K\trianglelefteq L\leq N_{G}(H)$, with $[K:H]=2$, $f(K)\equiv f(H)\mod 2$ if $L/K$ is cyclic of order 4 and $f(K)\equiv f(H)\mod 4$ is $L/K$ is quaternion of order 8.
  3. (c)
    For any elementary abelian subquotient $T/S$ of $G$ of rank 2, the equality
    $f(S)-f(T)=\sum_{S<X<T}\big{(}f(X)-f(T)\big{)}$
    holds.
The collection of Borel-Smith functions forms an additive subgroup $\operatorname{CF}_{b}(G,p)$ of $\operatorname{CF}(G,p)$. In fact, if $G$ is a finite $p$-group, then under the identification $\operatorname{CF}(G)\cong B^{*}(G):=\operatorname{Hom}(B(G),\mathbb{Z})$, $\operatorname{CF}_{b}(G)$ forms a rational $p$-biset subfunctor of $\operatorname{CF}(G)$. See [BY07, Proposition 3.7] for details.
Theorem 1.13.
We have a complete characterization of endotrivials.
  1. (a)
    [Mil25a, Theorem 4.6] Let $G$ be a finite $p$-group. The h-mark homomorphism is an isomorphism onto the subgroup of Borel-Smith functions
    $h\colon\operatorname{Pic}(\mathscr{K}(G))\cong\operatorname{CF}_{b}(G).$
  2. (b)
    [Mil25a, Corollary 6.4] Let $G$ be a finite group. The h-mark homomorphism has image $\operatorname{CF}_{b}(G,p)$ and induces a split exact sequence
    $0\to\operatorname{Hom}(G,k^{\times})\to\operatorname{Pic}(\mathscr{K}(G))\to\operatorname{CF}_{b}(G,p)\to 0.$
Example 1.14.
Let $G$ be a finite $p$-group. The first nontrivial example of an endotrivial, and those which make the magic happen in the elementary abelian case, are as follows (c.f. [BG25, Definition 12.3]). If $p=2$, one has an endotrivial for $C_{2}$ given by $kC_{2}\to k,$ with $k$ in homological degree 0 and the nonzero differential given by the augmentation homomorphism. Otherwise if $p$ is odd, one has an endotrivial for $C_{p}$ by truncating a periodic resolution of the trivial $kC_{p}$-module $k$, $kC_{p}\to kC_{p}\to k$ with $k$ in homological degree 0. For any subgroup $N\triangleleft G$ of index $p$, inflation yields the following complex
$u_{N}:=k[G/N]\to k[G/N]\to k$
Here $p$ is assumed odd, if $p=2$ then the highest degree term can be deleted.
A less easy example of an endotrivial for $p=2$ and $G=D_{2^{n}}$ is as follows. Let $H_{1},H_{2}$ be nonconjugate, noncentral subgroups of order 2 (this choice is unique up to conjugacy and reordering). Then the following complex is endotrivial, and a small example of a nontrivial faithful, irreducible endotrivial, i.e. one arising from a faithful irreducible real representation (see Definition 2.5),
$kD_{2^{n}}\to k[D_{2^{n}}/H_{1}]\oplus k[D_{2^{n}}/H_{1}]\to k,$
where all maps between indecomposable permutation modules are augmentation homomorphisms.
2. Borel-Smith functions and representation spheres
We begin by considering the topological properties of the main characters of our story, the endotrivials arising from genuine real representations, i.e. representations spheres. Such endotrivials come from objects of higher structure; they arise from chain complexes of free modules over the orbit category $\Gamma_{G}$, and in particular, their differentials are augmentation homomorphisms which are well-behaved with respect to local-global considerations. We discuss this now.
Definition 2.1.
We say that a superclass function $f$ is effective or if the following holds: if $K\leq H$ are subgroups of $G$, then $f(K)\geq f(H)$. That is, $f$ is monotonically decreasing with respect to the post of subgroups of $G$. Similarly, we say that an endotrivial $C$ is effective if its corresponding h-mark function $h_{C}$ is effective.
This choice of terminology will become clear in the sequel. We first recall some important facts about Borel-Smith functions. Let $G$ be a nilpotent group and $F$ a field. Given a $FG$-module $V$, the dimension function associated to $V$ is the superclass function
$\dim\colon H\mapsto\dim_{F}V^{H}.$
If $F$ has characteristic 0, this induces a group homomorphism $R_{F}(G)\to\operatorname{CF}(G)$. When $F=\mathbb{R}$, topologists may write $RO(G)=R_{\mathbb{R}}(G)$, as a representation sphere is nothing more than a one-point compactification of a real representation.
Theorem 2.2.
[tD87, Theorem 5.4 and Theorem 5.13, pages 211 and 216] Let $G$ be a nilpotent group. The image of $\dim\colon RO(G)\to\operatorname{CF}(G)$ is the group of Borel-Smith functions $\operatorname{CF}_{b}(G)$. Moreover, if $f$ is a effective Borel-Smith function, there exists a real representation $V$ for which $\dim(V)=f$.
Corollary 2.3.
Let $G$ be a nilpotent group and suppose $f$ is a Borel-Smith function. Then $f$ can be expressed as the difference of two effective Borel-Smith functions. In particular, every endotrivial is isomorphic in $\mathscr{K}(G)$ to the product of an effective endotrivial and the dual of an effective endotrivial.
Proof.
This follows from [tD87, Theorem 5.4, page 211] since the dimension function of any representation is effective, and [Mil25a, Theorem 4.6], since the image of the h-mark homomorphism is precisely the group of Borel-Smith functions of $G$. ∎
The previous corollary tells us that it suffices to work with effective endotrivials, which are particularly well-behaved. In fact, when $G$ is a nilpotent group, the dimension function gives a canonical effective basis of $\operatorname{CF}_{b}(G)$, which arises from the real irreps of $G$. Therefore, if $G$ is a finite $p$-group, we have a corresponding canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K})$ consisting of effective endotrivials.
Theorem 2.4.
Let $V_{1},\dots,V_{n}$ denote the irreducible real representations of $G$. Then the set of associated dimension functions $\{f_{1},\dots,f_{n}\}$, after removing duplicates, forms a $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$. In particular, there is an associated canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$.
Proof.
Since the image of the dimension homomorphism is precisely $\operatorname{CF}_{b}(G)$, it suffices to show that the set $\mathcal{B}:=\{f_{1},\dots,f_{n}\}$ is linearly independent after removing duplicates. [tD87, Proposition 5.9, page 213] asserts that $\ker(\dim)$ is generated by elements of the form $V-\psi^{k}(V)$, where $V$ is an irreducible real representation, $\psi^{k}$ is the $k$-th Adams operation, and $k$ is coprime to $|G|$. Therefore, the duplicates in $\mathcal{B}$ arise from Adams operation conjugates, and it follows that any set of real irreducible representations which are not Adams operation conjugates will correspond to a linearly independent set of Borel-Smith functions. ∎
Definition 2.5.
We call the Borel-Smith functions (resp. endotrivials) associated to the real irreps of $G$ the irreducible Borel-Smith functions (resp. endotrivials). These objects are necessarily effective.
Remark 2.6.
Given a real representation $V$ of $G$ with corresponding character $\chi$, we have an equality
$\dim_{\mathbb{R}}V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi(h),$
a practical method of computing the basis of $\operatorname{CF}_{b}(G)$.
Checking character tables of $G$ a finite $p$-group of normal $p$-rank one shows the basis of $\operatorname{Pic}(\mathscr{K}(G))$ obtained in [Mil24, Section 6] coincides with the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$. (The author remarks that he finds this is rather surprising, as the computations of [Mil24] were performed entirely ad-hoc and were done prior to the classification of $\operatorname{Pic}(\mathscr{K}(G))$.)
Notation 2.7.
Let $G$ be a finite $p$-group. Given a $\mathbb{R}G$-module $V$, let $f_{V}$ and $C_{V}$ denote the corresponding effective Borel-Smith function and effective indecomposable endotrivial respectively. That is, $f_{V}:=\dim V$ and $C_{V}$ is the unique indecomposable endotrivial for which $h_{C_{V}}=\dim V$.
Per tom Dieck [tD87], Borel-Smith functions correspond to representation spheres, which is simply a manifestation of 2.2. If $X$ is a finite-dimensional $G$-complex which is a $\mathbb{F}_{p}$-cohomology sphere, the fixed-point set $X^{H}$ is a $\mathbb{F}_{p}$-cohomology-sphere of dimension $d(H)$. The assignment $H\mapsto d(H)+1$ is (unfortunately) also called the dimension function of $X$. The set of dimension functions of representation spheres is precisely the effective non-negative Borel-Smith functions. Given a real representation $V$, let $S^{V}$ denote the corresponding representation sphere.
3. Chain complexes over the orbit category
Since effective endotrivials identify with representation spheres, they have additional structural properties, as representation spheres (and more generally, any $G$-CW-complex) produce free chain complexes over the orbit category $\Gamma_{G}$. We first review this notion, following [Yal17], then prove some technical results which will be critical for the sequel.
Definition 3.1.
Let $\Gamma_{G}$ denote the orbit category of $G$. The objects of $\Gamma_{G}$ are transitive $G$-sets $G/H$ for subgroups $H\leq G$, and the morphisms from $G/H$ to $G/K$ are $G$-set homomorphisms $G/H\to G/K$. In particular, $\operatorname{Hom}_{\Gamma_{G}}(H,K)$ is empty unless a conjugate of $H$ is a subgroup of $K$.
A $k\Gamma_{G}$-module $M$ is a contravariant functor from the category $\Gamma_{G}$ to the category of $k$-modules. By identifying $\operatorname{Aut}_{\Gamma_{G}}(G/H)$ with G//H{G}/\!\!/{H}, $M(H)$ has k[G//H]k[{G}/\!\!/{H}]-module structure. The category of finitely generated $k\Gamma_{G}$-modules, denoted $\operatorname{mod}(\Gamma_{G})$ is abelian, so the usual categorical concepts apply.
Given a $G$-set $X$, we define the $k\Gamma_{G}$-module $k[X^{?}]$ as the module with value at $G/H$ given by $k[X^{H}]$, with the obvious induced maps. A module over $\Gamma_{G}$ is free if it is isomorphic to a direct sum of modules of the form $k[(G/K)^{?}]$. Let $\operatorname{proj}(\Gamma_{G})$ denote the full subcategory of $\operatorname{mod}(\Gamma_{G})$ consisting of free $\Gamma_{G}$-modules (by the Yoneda lemma, every projective $\Gamma_{G}$-module is free, see [Yal17, Definition 3.1]).
The next propositions follow immediately from definition of $k\Gamma_{G}$-modules. We refer to these facts as the stabilizers grow conditions. This property and name were first suggested to the author by Robert Boltje.
Proposition 3.2.
Let $X$ and $Y$ be two $G$-sets. Any $k\Gamma_{G}$-homomorphism $f\colon k[X^{?}]\to k[Y^{?}]$ satisfies the following property: given any $x\in X$ and subgroup $H\leq G$, if $x\in X^{H}$, then $f(x)\in k[Y^{H}]$. In particular, $\operatorname{Hom}_{\operatorname{mod}(\Gamma_{G})}(k[X^{?}],k[Y^{?}])$ has $k$-basis induced from the set $\operatorname{Hom}_{G-\operatorname{set}}(X,Y)$.
Proposition 3.3.
Let $k[X^{?}]$ and $k[Y^{?}]$ be two free $k\Gamma_{G}$-modules and let $f\colon k[X^{?}]\to k[Y^{?}]$ be a $k\Gamma_{G}$-module homomorphism. Let $K\leq_{G}H$ be two subgroups of $G$. For any $m\in k[X^{H}]$, if $f^{H}(m)=n\in k[Y^{H}]$, then regarding $m$ as an element of $k[X^{K}]\supseteq k[X^{H}]$ of k[G//H]k[{G}/\!\!/{H}]-modules, $f^{K}(m)=n\in k[Y^{K}]$.
Remark 3.4.
As a result, each free $k\Gamma_{G}$-module has a stratification by fixed points. That is, given a free $k\Gamma_{G}$-module $kX$, we may write $kX=kG^{\oplus a_{1}}\oplus k[G/H_{2}]^{\oplus a_{2}}\cdots\oplus k[G/G]^{a_{n}}$ for some $a_{1},\dots,a_{n}\in\mathbb{N}$, where each direct summand $k[G/H_{i}]^{\oplus a_{i}}$ is a uniquely determined submodule of $kX$. Therefore, we may fix a “canonical permutation basis” in accordance with this stratification. We will make considerable use of this fact in the sequel.
Given a $G$-set $X$ and two subgroups $K\leq H$ of $G$, we have a canonical inclusion homomorphism (of $k[N_{G}(H)\cap N_{G}(K)]$-modules) $i\colon k[X^{H}]\to k[X^{K}]$ associated to the $G$-set homomorphism $G/K\to G/H,1K\mapsto 1H$. This comes associated with a (not necessarily unique) projection map $p\colon k[X^{K}]\to k[X^{H}]$ satisfying $p\circ i=\operatorname{id}$ and $p(m)=0$ if $m\in\Psi^{H}(k[X^{K}])=0$. After choosing a canonical permutation basis, we obtain a corresponding projection map associated to the basis.
For any subgroup $H\leq G$, we have an obvious functor $(H)\colon\operatorname{proj}(\Gamma_{G})\to\operatorname{perm}(k[{G}/\!\!/{H}])$ from the assignment $k[X^{?}]\to k[X^{H}]$. Modular fixed points behave as one would hope.
Proposition 3.5.
Let $P$ be a $p$-subgroup of $G$. Then the following diagram commutes up to natural isomorphism.
$\operatorname{proj}(\Gamma_{G})$$\operatorname{perm}(kG)$$\operatorname{perm}(k[{G}/\!\!/{P}])$$\scriptstyle{-(1)}$$\scriptstyle{-(P)}$$\scriptstyle{\Psi^{P}}$
Proof.
This follows from the natural isomorphism $k[X^{P}]\cong\Psi^{P}(kX)$. ∎
Definition 3.6.
Let $X$ be a $G$-CW-complex. The reduced chain complex $X$ over the orbit category is the functor $\tilde{C}_{\ast}(X^{?};k)$ from the orbit category $\Gamma_{G}$ to the category of chain complexes of $k$-modules. This gives rise to a chain complex of free (by [Yal17, Lemma 3.2]) $k\Gamma_{G}$-modules
$\tilde{C}_{\ast}(X^{?};k):=\cdots\to C_{i}(X^{?};k)\xrightarrow{d_{i}}C_{i-1}(X^{?};k)\to\cdots\to C_{0}(X^{?};k)\xrightarrow{\epsilon}\underline{k}\to 0,$
where $\underline{k}$ denotes the constant functor with values $k(H)=k$ for all subgroups $H\leq G$, and $\epsilon$ denotes the augmentation homomorphism (this is essentially the construction of reduced Bredon homology, see [Ill73]). We denote chain complexes of free $k\Gamma_{G}$-modules by $C^{?}$, and their evaluations at the subgroup $H\leq G$ by $C^{H}$ to inspire an aura of fixed points.
If $G$ is a finite $p$-group and $S$ is a representation sphere of $G$, it may be realized as a $G$-CW-complex, hence producing a chain complex of free $k\Gamma_{G}$-modules. Therefore for any subgroup $H\leq G$, $\tilde{C}_{\ast}(S^{H};k)$ is an effective endotrivial complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules. Conversely, after shifting, every effective endotrivial complex $C$ of $kG$-modules is homotopy equivalent to $\tilde{C}_{\ast}(S^{1};k)$ for some representation sphere $S$ of $G$ with dimension homomorphism $\dim(S)=h_{C}$ (see [tD87, Page 217, Theorem 5.16]), and hence up to homotopy arises from a chain complex of free $k\Gamma_{G}$-modules.
This proposition is critical for many of the constructions to follow: it allows us to perform lifts of $p$-local homomorphisms $k\to\Psi^{H}(C)$.
Proposition 3.7.
Let $G$ be a finite $p$-group, let $k[X^{?}]$ and $k[Y^{?}]$ be two free $k\Gamma_{G}$-modules with $X$ transitive, let $f\colon k[X^{?}]\to k[Y^{?}]$ be a $k\Gamma_{G}$-module homomorphism, and let $H\leq G$ be a subgroup. Suppose $k[X^{H}]\neq 0$, and let $M^{\prime}$ denote the unique minimal nonzero submodule of $k[X^{H}]$. If $M^{\prime}\subseteq\ker(f^{H})$, then there exists a submodule $M^{?}\subseteq k[X^{?}]$, minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$, such that $M^{H}=M^{\prime}$ and for all subgroups $K\leq_{G}H$, $M^{K}\subseteq\ker(f^{K})$.
Proof.
We define $M^{?}\subseteq k[X^{?}]$ as follows: for $L\leq G$, $M^{L}\subset k[X^{L}]$ is the unique minimal submodule if $X$ has any $K$-fixed points, and the zero module if not. It is straightforward that, by construction, $M^{?}$ is minimal with respect to the property that $M^{L}\neq 0$ when $k[X^{L}]\neq 0$. Now by minimality of $M^{?}$, it suffices to show $f^{K}$ is not an injective k[G//K]k[{G}/\!\!/{K}]-module homomorphism for all $K\leq_{G}H$, but this follows directly from Proposition 3.3. ∎
Remark 3.8.
Proposition 3.7 asserts that we can construct “local” inclusions $k\to C[s]$ for chain complexes arising from chain complexes over the orbit category. This does not hold for arbitrary objects of $\mathscr{K}(G)$. For instance, the two-term complex $C=k\xrightarrow{\operatorname{coaug}}kG$ (with $k$ in homological degree 0) has no nonzero global homomorphism $k\to C$, but there exist nonzero local homomorphisms $k\to\Psi^{H}(C)$ (in fact isomorphisms) for every nontrivial subgroup $H\leq G$, since $\Psi^{H}(kG)=0$.
3.1. Removing contractible summands
We prove a couple technical results for finite $p$-groups that show we can remove contractible summands from chain complexes of free $k\Gamma_{G}$-modules if and only if we can for the corresponding chain complexes of permutation $kG$-modules (in particular, for effective endotrivials). Though most of these results are not strictly necessary for future sections, they provide elucidation and may be of independent interest.
Remark 3.9.
A chain complex of free $k\Gamma_{G}$ modules has in each homological degree a canonical permutation basis, and each differential respects the stabilizers grow condition. Moreover, it is easy to see that given a chain complex $C$ of permutation $kG$-modules, if one can choose a canonical permutation basis of $G$-sets in each degree such that the stabilizers grow condition holds with respect to this basis, then $C$ may be realized as the image of a chain complex of free $k\Gamma_{G}$-modules evaluated at the subgroup $1\leq G$. In this sense, the chain complexes of permutation $kG$-modules satisfying the stabilizers grow condition on differentials are exactly those which arise from chain complexes of free $k\Gamma_{G}$-modules.
If $C$ is a chain complex of permutation modules, we say $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules if there exists a chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C^{1}\cong C$ as chain complexes.
Proposition 3.10.
Let $G$ be a finite group and suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. Let $i$ be an integer and suppose $T\subseteq C_{i}$ is a $k$-dimension one submodule. If $T\not\in\ker(d_{i})$, then there exists a contractible direct summand $K$ of $C$ consisting of permutation modules such that $T\subseteq K_{i}$.
Proof.
We have a canonical permutation basis associated to $C_{i}$ arising from the $k\Gamma_{G}$-module structure, $C_{i}\cong kX_{1}\oplus\cdots\oplus kX_{m}$ with each $X_{j}$ a transitive $G$-set. For $j\in\{1,\dots,m\}$, let $T_{j}$ denote the projection of $T$ into $kX_{j}$. Similarly, we have a canonical permutation basis $C_{i-1}\cong kY_{1}\oplus\cdots\oplus kY_{n}$ with each $Y_{j}$ a transitive $G$-set. For $j\in\{1,\dots,n\}$, let $U_{j}$ denote the projection of $T$ into $kY_{j}$. Finally, let $\{b_{1},\dots,b_{l}\}$ denote the indices for which the projection $p_{j}\circ d_{i}(T)$ onto $U_{j}$ is nonzero. By assumption this set is nonempty. Set $b$ equal to the index $b_{j}$ for which $Y_{b_{j}}$ has minimal stabilizers (this is well-defined since $Y_{b_{j}}$ is transitive).
It follows from Corollary 3.3 that there exists an index $a\in\{1,\dots,m\}$ such that the composition $p_{b}\circ d_{i}\circ i_{a}$ is an isomorphism, where $p_{b}$ denotes projection onto $kY_{b}$ and $i_{a}$ denotes inclusion into $kX_{a}$. In particular, $X_{a}$ and $Y_{b}$ are isomorphic $G$-sets. Note the choice of $a$ is not necessarily unique (for example, the two-term complex $k\oplus k\to k$ with differential $(\operatorname{id},\operatorname{id})$).
Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{a}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}^{\prime}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{b}\oplus\cdots\oplus kY_{n}$. We will now modify the canonical bases of $C_{i}$ and $C_{i-1}$ to construct the contractible chain complex $K$ that splits off of $C$. Choose any generator $t\in T$ of $T$, then we may write $t=p_{a}(t)+t^{\prime}$, with $t^{\prime}\in T_{1}\oplus\cdots\oplus\widehat{T}_{a}\oplus\cdots\oplus T_{m}$. The $kG$-modules $\langle p_{a}(t)\rangle$ and $\langle t^{\prime}\rangle$ are both isomorphic to $k$, i.e. are $G$-stable. We replace the $C_{i}$ basis elements $\{x_{1},\dots,x_{e}\}=X_{a}\subset kX_{a}\subseteq C_{i}$ with $\{x_{1}+t^{\prime},\dots,x_{e}+t^{\prime}\}=X_{a}^{\prime}\subset C_{i}$. $X_{a}^{\prime}$ is again a $G$-set, and we have $C_{i}=C_{i}^{\prime}\oplus kX^{\prime}_{a}$. Moreover, $T\subseteq kX_{a}^{\prime}$ by construction.
We replace the $C_{i-1}$ basis elements $\{y_{1},\dots,y_{f}\}=Y_{b}\subset kY_{b}\subseteq C_{i-1}$ with $\{d_{i}(x_{1}+t^{\prime}),\dots,d_{i}(x_{e}+t^{\prime})\}=X_{b}^{\prime}\subset C_{i-1}$. Again, $Y_{b}^{\prime}$ is a $G$-set. We claim that $p_{b}(X_{a}^{\prime})$ is a $k$-basis of $Y_{b}$. This follows because $\langle p_{b}\circ d_{i}(t^{\prime})\rangle$ is either 0 or the unique $k$-dimension one submodule of $kY_{b}$, and we chose $b$ such that both $p_{b}\circ d_{i}\circ i_{a}$ was an isomorphism and $p_{b}\circ d_{i}(T)$ was nonzero. Therefore $p_{b}\circ d_{i}(y_{j}+t^{\prime})$ has the same $G$-stabilizer as both $y_{j}$ and $y_{j}+t^{\prime}$. Hence $Y_{b}^{\prime}$ is isomorphic as a $G$-set to $Y_{b}$ and $X_{a}$, and it follows that $C_{i-1}=C^{\prime}_{i-1}\oplus kY^{\prime}_{b}$. Moreover, $d_{i}$ restricts to an isomorphism $kX_{a}^{\prime}\xrightarrow{\cong}kY_{b}^{\prime}$, and the result follows. ∎
The next proposition demonstrates that if a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules satisfies that $C^{1}$ has a contractible summand (as a chain complex of $kG$-modules), then $C^{?}$ itself has a contractible summand (as a chain complex of $k\Gamma_{G}$-modules).
Proposition 3.11.
Let $G$ be a finite group and suppose $C$ is a bounded chain complex of permutation $kG$-modules arising from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. If $C$ contains a contractible direct summand $K$, then there exists a direct sum decomposition $C=K\oplus D$ such that $D$ also arises from a chain complex $D^{?}$ of free $k\Gamma_{G}$-modules.
In particular, $C$ has a contractible direct summand of permutation modules if and only if $C^{?}$ has a contractible direct summand.
Proof.
To prove this, it suffices to find a decomposition of $C$ into direct summands $K,D$ as stated and show that each homological degree of $D$ has a canonical permutation basis satisfying the stabilizers grow condition. First, since $C$ has a contractible direct summand, there exists an integer $i\in\mathbb{Z}$ and $k$-dimension one submodule $T\subseteq C_{i}$ such that $T\not\in\ker(d_{i})$. Let $C_{i}=kX_{1}\oplus\cdots\oplus kX_{m}$ and $C_{i-1}=kY_{1}\oplus\cdots\oplus kY_{n}$ be the direct sum decompositions into canonical bases, with each $X_{j},Y_{j}$ a transitive permutation module. After projecting $T$ onto each $kX_{j}$, it follows that there exists an index $l$ for which $d_{i}|_{X_{l}}$ is injective. Choose the index $l$ such that the stabilizer of $X_{l}$ is minimal with respect to subgroup inclusion.
We modify the canonical bases of $C_{i}$ and $C_{i-1}$ as follows. First, it follows from Corollary 3.3 that there exists a (non-unique) index $h\in\{1,\dots,n\}$ such that $p_{h}\circ d_{i}\circ i_{l}$ is an isomorphism, where $i_{l}$ denotes inclusion into $kX_{l}$ and $p_{h}$ denotes projection onto $kY_{h}$. We replace the basis elements of $Y_{h}$ with the basis elements $Y_{h}^{\prime}=\{d_{i}(x)\mid x\in X_{l}\}$. Then $Y_{h}^{\prime}$ is a $G$-set isomorphic to $X_{l}$, and since the projection of $kY_{h}^{\prime}$ onto $kY_{h}$ is an isomorphism, this forms a new basis of $C_{i-1}$. Finally, we have $d_{i-1}(kY_{h}^{\prime})=0$ since $d_{i-1}\circ d_{i}=0$, so the resulting basis still respects the stabilizers grow property. Set $D_{i-1}:=kY_{1}\oplus\cdots\oplus\widehat{kY}_{h}\oplus\cdots\oplus kY_{n}$ and $K_{i-1}:=kY_{h}^{\prime}$.
Set $K_{i}:=kX_{l}$, by construction $d_{i}|_{K_{i}}$ is an isomorphism and maps canonical permutation basis to canonical permutation basis. We next modify the canonical permutation basis of $C_{n}$ to obtain a direct summand $D_{i}$ for which $d_{i}(D_{i})\subseteq D_{i-1}$. Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{l}\oplus\cdots\oplus kX_{m}$, so we have $C_{i}=C_{i}^{\prime}\oplus K_{i}$. For each canonical permutation basis element $x$ of $C_{i}$ (i.e. some $x\in X_{j}$ for $j\neq l$) replace $x$ with $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$, where $p_{h}^{\prime}$ denotes projection onto $kY_{h}^{\prime}$ and $(d_{i}|_{kX_{l}})^{-1}$ denotes the inverse of the isomorphic projection of $kX_{l}$ onto $kY_{h}^{\prime}$. A straightforward computation shows $d_{i}(x^{\prime})\in D_{i-1}$. By construction, $((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)\in K_{i}$, and the collection $X^{\prime}$ of the $x^{\prime}$ forms a $G$-set, with $X^{\prime}\cong X_{1}\sqcup\cdots\sqcup\widehat{X}_{l}\sqcup\cdots\sqcup X_{m}$. Set $D_{i}:=kX^{\prime}$, then we have a direct sum decomposition $C_{i}=D_{i}\oplus K_{i}$. Finally, for all $j\neq i,i-1$, set $D_{j}=C_{j}$. It follows that we have a direct sum decomposition $C=K\oplus D$. Finally, observe that for each $x^{\prime}$, under the image of $d_{i}$, the canonical permutation basis representation of $x^{\prime}$ in $D_{i-1}$ is identical to the canonical permutation basis representation of $x$ in $D_{i-1}$, therefore the stabilizers grow property holds for the differential $d_{i}$.
It remains to show the stabilizers grow property holds for the differentials $d_{i-1}$ and $d_{i+1}$ with respect to the same canonical bases for $C_{i+1}$ and $C_{i-1}$. The condition holding for $d_{i-1}$ is straightforward, since the only modified basis elements in $C_{i-1}$ now belong to $\ker(d_{i-1})$. Similarly, $\operatorname{im}(d_{i+1})\subseteq D_{i}$, and under the decomposition $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$ into $C_{i}^{\prime}$ and $kY^{\prime}_{h}$ respectively, it follows by construction that if a canonical permutation basis element $z\in C_{i+1}$ satisfies $d_{i+1}(z)=\sum a_{j}x_{j}+m$ with $m\in kX_{l}$ and each $x_{j}$ a canonical permutation basis element of $C_{i}^{\prime}$, then $d_{i+1}(z)=\sum a_{j}x_{j}^{\prime}$ where $x_{j}^{\prime}$ is the refined basis element of $X^{\prime}$ corresponding to $x_{j}$. Thus $d_{i+1}$ also satisfies the stabilizers grow condition, as desired.
In particular, if $G$ is a finite $p$-group, then every indecomposable effective endotrivial arises from a free chain complex of $k\Gamma_{G}$-modules.
Corollary 3.12.
Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial of $kG$-modules. Then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
Proof.
Recall every effective endotrivial (after a possible shift) corresponds to a real representation $V$ of $\mathbb{R}G$-modules. The representation sphere $S^{V}$ is a $G$-CW-complex, therefore produces a (reduced) chain complex $C^{?}$ for which $C^{1}$ is an endotrivial with the same h-marks as $C$. By Corollary 3.11, we can remove all contractible summands from $C^{?}$ until $C^{1}$ is indecomposable, and since two indecomposable endotrivials with the same h-marks are isomorphic, the result follows. ∎
Moreover, we can completely identify the homomorphisms $k\to C$ for an effective endotrivial $C$.
Corollary 3.13.
Let $G$ be a finite $p$-group and let $C^{?}$ be a chain complex of free $k\Gamma_{G}$-modules with no contractible summands (e.g. an indecomposable effective endotrivial). For every integer $s\in\mathbb{Z}$ and submodule $T\subseteq C^{1}_{i}$ isomorphic to $k$, there exist a chain complex homomorphism $f\colon k\to C[s]$ with $\operatorname{im}(f_{0})=T$. In particular, $\dim_{k}\operatorname{Hom}_{\mathscr{K}(G)}(k[0],C^{1}[s])$ is equal to the number of indecomposable direct summands of $C_{s}$.
Proof.
The existence of such a homomorphism $f$ is equivalent to the inclusion $T\subseteq\ker(d_{i}),$ and this inclusion holds since if not, Proposition 3.10 implies the existence of a contractible summand of $C^{1}$, hence a contractible summand of $C^{?}$, which cannot occur. The last statement is straightforward. ∎
4. An open cover of the Balmer spectrum via endotrivials
For the rest of the paper, we assume $G$ is a finite $p$-group.
4.1. $p$-local quasi-isomorphisms
Proposition 4.1.
Let $G$ be a finite $p$-group and let $V$ be a $\mathbb{R}G$-module with kernel $N\trianglelefteq G$. Then the associated descending indecomposable endotrivial $C_{V}$ satisfies the following property: $n:=h_{C_{V}}(1)$ is the maximal nonzero homological degree of $C_{V}$, and $(C_{V})_{n}\cong k[G/N]$.
Proof.
Set $C:=C_{V}$. First, $V$ is equivalently a faithful $\mathbb{R}[G/N]$-module, so it suffices to assume $N=1$ after replacing $G$ with $G/N$, and it suffices to show $C_{n}$ is indecomposable and projective. Since $C$ is effective, it follows by an inductive argument that for all subgroups $H\leq G$, $\Psi^{H}(C)$ is isomorphic to an indecomposable complex in $\mathscr{K}({G}/\!\!/{H})$ whose highest nonzero homological degree is $h_{C}(H)$ (see [Lin18, Proposition 5.8.11]). In particular, $n=h_{C}(1)$ is the maximal nonzero degree of $C$, since $C$ is indecomposable.
First, we show that there cannot exist any non-projective direct summands of $C_{n}$, and that more generally, the highest homological degree $i$ for which the permutation module $k[G/H]$ can occur as a direct summand of $C_{i}$ is $i=h_{C}(H).$ Since $V$ is a faithful representation, $h_{C}=\dim V$ satisfies $h_{C}(H)<h_{C}(1)$ for any nontrivial subgroup $H\leq G$. Let $j$ be the highest degree for which $\Psi^{H}(C)_{j}\neq 0$. If $j\leq h_{C}(H)$, there is nothing to show, as this implies that only permutation modules with stabilizers not contained in $H$ occur in degrees above $h_{C}(H)$. Otherwise, if $j>h_{C}(H)$, since $\Psi^{H}(C)$ is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to an indecomposable complex with highest nonzero degree $h_{C}(H)$, the differential $\Psi^{H}(C)_{j}\xrightarrow{\Psi^{H}(d_{j})}\Psi^{H}(C)_{j-1}$ is split injective. But now, we may write $C_{j}\xrightarrow{d_{j}}C_{j-1}$ as follows:
$M_{1}$$N_{1}$$M_{2}$$N_{2}$$\scriptstyle{d_{1}^{1}}$
Here, we have $C_{j}=M_{1}\oplus M_{2}$, $C_{j-1}=N_{1}\oplus N_{2}$, where $M_{1}$ and $N_{1}$ satisfy $\Psi^{H}(M_{1})=\Psi^{H}(C_{j})$ and $\Psi^{H}(N_{1})=\Psi^{H}(C)_{j-1}$ and $M_{2}$ and $N_{2}$ satisfy $\Psi^{H}(M_{2})=0$ and $\Psi^{H}(N_{2})=0$. With this setup, we have $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$. By applying [Mil24, Lemma 5.7] and its dual statement, $d_{1}^{1}$ is an isomorphism if and only if $\Psi^{K}(d_{1}^{1})$ is an isomorphism for all $K\geq H$. Since $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$ is an isomorphism, it follows by an inductive argument up the poset of subgroups $K$ of $G$ containing $H$ that $\Psi^{H}(d_{1}^{1})$ is an isomorphism. Now by a standard homological algebra argument (c.f. [BM23, Lemma 9.2]), $M_{1}\xrightarrow{d_{1}^{1}}N_{1}$ splits off as a contractible direct summand of $C$, a contradiction since $C$ was assumed to be indecomposable. Therefore, $j=h_{C}(H)$. We conclude the highest homological degree $i$ for which the permutation module $k[G/H]$ can occur as a direct summand of $C_{i}$ is $i=h_{C}(H).$ In particular, since $V$ is faithful, only projective modules can occur in $C_{n}$.
It remains to show $C_{n}$ is indecomposable. Suppose for contradiction $C_{n}\cong P_{1}\oplus P_{2}$ for projective $kG$-modules $P_{1},P_{2}$. Since $\dim_{k}H_{n}(C)=1$, either $d_{n}|_{P_{1}}$ or $d_{n}|_{P_{2}}$ is injective, hence split injective, and it follows that a contractible chain complex containing $P_{1}$ or $P_{2}$ splits off from $C$, contradicting indecomposability of $C$. Thus, $C_{n}$ is indecomposable projective, as desired. ∎
In the situation described in Proposition 4.1, $\ker(d_{n})$ is the unique submodule of $k[G/N]$ of $k$-dimension 1. We are now ready to generalize the $a_{N}$ and $b_{N}$ homomorphisms constructed in [BG25, Definition 12.3] for every effective endotrivial $C$ and ($p$-)subgroup $H\leq G$.
Theorem 4.2.
Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial. For every subgroup $H\leq G$, there exists a chain complex homomorphism
$\iota^{H}_{C}\colon k[h_{C}(H)]\to C$
unique up to scaling, such that $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism $k[h_{C}(H)]\cong\Psi^{H}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$. We have $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$. Moreover, the image of $\iota^{H}_{C}$ in $C_{h_{C}(H)}$ is contained in an indecomposable direct summand isomorphic to $k[G/K]$, for some subgroup $K\geq H$.
Proof.
Since $C$ is effective, we may assume (after possibly shifting $C$) by Corollary 3.12 that there exists an indecomposable chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C\cong C^{1}$. Denote the differentials of $C^{?}$ by $d_{i}^{?}$. By Proposition 4.1, the chain complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules $C^{H}$, which by Proposition 3.5 is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to $\Psi^{H}(C)$, contains up to homotopy a unique indecomposable k[G//H]k[{G}/\!\!/{H}]-module $M$ in top homological degree $h_{C}(H)$, and $\dim_{k}\ker(d_{h_{C}(H)}^{H})=1$. It follows by the stabilizers grow condition that in the canonical permutation basis of $C^{?}$, $M$ corresponds to a unique direct summand in homological degree $h_{C}(H)$.
Therefore, we are in the situation of Proposition 3.7 after restricting $d_{i}^{?}$ to $M$. Choosing a nonzero $m\in\ker(d_{h_{C}(H)}^{H})$ (this choice is unique up to scaling), Proposition 3.7 asserts the existence of a $m^{\prime}\in\ker(d_{h_{C}(H)}^{1})$ generating a (unique) $k$-dimension one submodule of $C^{1}_{h_{C}(H)}$. We have a unique (up to scaling) nonzero homomorphism $\iota^{H}_{C}$ of chain complexes as follows.
[…]
We have that $\Psi^{H}(\iota^{H}_{C})$ is by construction a quasi-isomorphism. The statement $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$ follows since $c_{g}\circ\Psi^{H}=\Psi^{H^{\prime}}$ for any $g\in G$ satisfying $^{g}H=H^{\prime}$. The final statement follows from the final statement of Proposition 3.7. ∎
Remark 4.3.
One has to take care in how the maps $\iota^{H}_{C}$ are constructed; such maps that become quasi-isomorphisms locally are non-unique (not even up to choice of identification as in [BG25, Remark 12.5]). For instance, let $G=D_{16}$, $C$ be the endotrivial
$C:=kD_{16}\xrightarrow{d_{2}}k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}]\xrightarrow{d_{1}}k,$
where $H_{1}$ and $H_{2}$ are non-conjugate non-central subgroups of order 2, $d_{2}$ is induced by the assignment $g\mapsto(gH_{1},gH_{2})$, and $d_{1}$ is induced by augmentation homomorphisms. In this case, we could choose $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ to be the inclusion
$\iota\colon k\to k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}],\quad 1\mapsto\left(\sum_{g\in[G/H_{1}]}gH_{1},\sum_{g\in[G/H_{2}]}gH_{2}\right).$
In this case, both $\hat{\Psi}^{H_{1}}(\iota)$ and $\hat{\Psi}^{H_{2}}(\iota)$ are isomorphisms, a desired property. However, from the construction in Proposition 3.7, it follows that $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ satisfy $\Psi^{H_{1}}(\iota^{H_{2}}_{C})=0$ and vice versa, since neither $H_{1}\geq_{G}H_{2}$ or vice versa.
Additionally, given two subgroups $K,H\leq G$, $\iota^{H}_{C}$ and $\iota^{K}_{C}$ may coincide. For instance, any shift of the tensor unit has $\iota^{H}_{C}=\iota^{G}_{C}$ for all subgroups $H\leq G$.
Proposition 4.4.
Let $H,K$ be subgroups of $G$ and let $C$ be an effective endotrivial of $kG$-modules. Then $\iota^{H}_{C}=\iota^{K}_{C}$ if and only if there exists a subgroup $B$ of $G$ such that $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$.
Proof.
The forward implication is straightforward by the construction in Proposition 4.2 - in particular, we may choose $B$ to be the vertex of the permutation $kG$-submodule containing the image of $\iota^{H}_{C}$. Conversely, assume $h_{C}(H)=h_{C}(K)$. Suppose there exists a subgroup $B$ satisfying $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$. From the construction in Proposition 4.2, it follows that the image of $\iota^{B}_{C}$ is the unique minimal submodule of a transitive permutation module isomorphic to $k[G/B^{\prime}]$ for some $B^{\prime}$ containing $B$. It follows that $\iota^{H}_{C}$ and $\iota^{K}_{C}$ also have image contained in this permutation module, and hence also have image the unique minimal submodule of $k[G/B^{\prime}]$, as desired. ∎
4.2. The open cover
Construction 4.5.
With the notation of [BG25, Definition 12.3], given any normal subgroup $N$ of $G$ of index $p$, the maps $a_{N}$ and $b_{N}$ are examples of such morphisms (however, $c_{N}$ is not!). For instance, the endotrivial
$u_{N}=k[G/N]\to k[G/N]\to k$
(assuming $p$ odd) satisfies $a_{N}=\iota_{u_{N}}^{H}$ for any $H\not\leq N$ and $b_{N}=\iota_{u_{N}}^{H}$ for any $H\leq N$.
Given this observation, we generalize the open cover of $\operatorname{Spc}(\mathscr{K})$ presented in [BG25, Proposition 13.11] as follows. Let $\mathcal{B}(G)$ denote the subset of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ not induced from the trivial $\mathbb{R}G$-module $\mathbb{R}$, i.e. the effective endotrivials arising from irreducible real representation spheres, excluding $k[1]$. Define an open of $\operatorname{Spc}(\mathscr{K})$ by
$U(H):=\bigcap_{C\in\mathcal{B}(G)}\operatorname{open}(\iota_{C}^{H}).$
Here,
$\operatorname{open}(f):=\operatorname{open}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is invertible in }\mathscr{K}(G)/\mathscr{P}\}.$
Note that in defining the opens $U(H)$, the canonical $\mathbb{Z}$-basis element $k[1]\in\mathcal{B}(G)$ would play no significant role, so we may exclude $k[1]$ from $\mathcal{B}(G)$ without issue.
Remark 4.6.
Let us recall the closed points of $\operatorname{Spc}(\mathscr{K})$. Recall that closed points are exactly the minimal primes for inclusion, and every prime contains a minimal prime. [BG25, Corollary 7.31] asserts that the minimal primes of $\operatorname{Spc}(\mathscr{K})$ are those explicitly of the form $\mathfrak{m}_{H}:=\mathscr{P}(H,0)$ for some subgroup $H$, where 0 denotes the unique closed point of $V_{{G}/\!\!/{H}}$. These closed points are precisely the kernel of the residue tt-functor $\mathbb{F}^{H}:=\operatorname{Res}_{1}^{{G}/\!\!/{H}}\circ\Psi^{H}=\Psi^{H}\circ\operatorname{Res}^{G}_{H}$ (see [BG25, Definition 7.26]).
The collection $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, with the closed point $\mathfrak{m}_{H}$ belonging to $U(H)$.
Proposition 4.7.
Let $H$ be a subgroup of $G$ and $C$ be an effective endotrivial. Recall the residue tt-functor $\mathbb{F}^{H}\colon\operatorname{Res}^{{G}/\!\!/{H}}_{1}\circ\Psi^{H}\colon\mathscr{K}\to\operatorname{D}_{b}(k)$ of the closed point $\mathfrak{m}_{H}$. $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism. Moreover, if $K$ is a subgroup of $G$ satisfying $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism.
In particular, the open $U(H)$ contains $\mathfrak{m}_{H}$, therefore the set of opens $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$.
Proof.
The fact that $\mathbb{F}^{H}(\iota^{H}_{C})$ is an isomorphism follows immediately since $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism. Moreover, if $h_{C}(K)\neq h_{C}(H)$, $\mathbb{F}^{K}(\iota^{H}_{C})$ cannot be an isomorphism since $k[h_{C}(H)]$ is not isomorphic to $\Psi^{K}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{K}])$.
It follows by conservativity that for any effective endotrivial $C$, $\iota^{H}_{C}$ is an isomorphism in $\mathscr{K}/\mathfrak{m}_{H}$ since $\mathbb{F}^{H}$ is the residue functor of $\mathfrak{m}_{H}$, so $U(H)$ contains $\mathfrak{m}_{H}$. Therefore by general tt-geometry, $\{U(H)\}_{H\leq G}$ is an open cover of $\operatorname{Spc}(\mathscr{K})$, as every prime specializes to some $\mathfrak{m}_{H}$, which are precisely the closed points of $\operatorname{Spc}(\mathscr{K}(G)$. ∎
We next show that if $K,H$ are non-conjugate subgroups of $G$, then $\mathfrak{m}_{K}\not\in U(H)$. This takes a bit more work. It follows from the previous proposition that if there exists a Borel-Smith function $f$ for which $f(K)\neq f(H)$, then $\mathfrak{m}_{K}\not\in U(H)$. However, this may not necessarily occur!
Definition 4.8.
Let $K,H$ be a pair of non-conjugate subgroups of $G$. Say $K$ and $H$ are indistinguishable if for all Borel-Smith functions $f$, $f(K)=f(H)$. If no such indistinguishable pairs exist in $G$, say non-conjugacy is detected in $G$.
Remark 4.9.
If $K$ and $H$ are indistinguishable, we have $|K|=|H|$, since $\dim(\mathbb{R}G)^{H}=[G:H]$, the index of $H$ in $G$.
Many $p$-groups have non-conjugacy detected, such as abelian groups trivially, all finite $p$-groups of normal $p$-rank one (this is computed indirectly in [Mil24, Section 6]), and all groups of order at most $p^{3}$. However, indistinguishable pairs of subgroups exist. For instance, set $G:=C_{8}\ltimes(C_{2}\times C_{2})$, with generators $a,b,c$ satisfying
$a^{8}=b^{2}=c^{2}=1,\quad{}^{b}a=a^{-1},\quad{}^{c}a=a^{3},\quad bc=cb.$
This group has a GAP implementation of SmallGroup(32,43); it is the holomorph of $\mathbb{Z}/8\mathbb{Z}$. In this case, $C_{2}\times C_{2}$ acts on $D_{8}$ faithfully. $G$ has two nonconjugate subgroups isomorphic to $V_{4}$,
$H:=\{1,b,c,bc\},\quad K:=\{1,a^{2}b,a^{2}c,bc\},$
whose individual elements are all conjugate:
$b\sim_{G}a^{2}b,\quad c\sim_{G}a^{2}c.$
Therefore, given any real representation $V$ of $G$ with character $\mathscr{P}_{V}$, we have $\dim V^{H}=\dim V^{K}$, since
$\dim V^{H}=\frac{1}{|H|}\sum_{h\in H}\chi_{V}(h).$
Since $\operatorname{im}(\dim)=\operatorname{CF}_{b}(G)$, every Borel-Smith function $f$ satisfies $f(H)=f(K)$. The author thanks Math.SE user testaccount for this example.
Lemma 4.10.
Let $H,K$ be non-conjugate subgroups of $G$ with the same order, let $L$ be the smallest subgroup of $G$ containing both a conjugate of $H$ and a conjugate of $K$, and let $C$ be an effective endotrivial for which $h_{C}(L)<h_{C}(H)$. Then $\mathbb{F}^{K}(\iota^{H}_{C})$ and $\mathbb{F}^{H}(\iota^{K}_{C})$ are 0.
Proof.
It suffices to assume $h_{C}(H)=h_{C}(K)$. From Proposition 4.1 and the construction of $\iota^{H}_{C}$ in Theorem 4.2, the image of $\iota^{H}_{C}$ in homological degree $h_{C}(H)$ is contained in an indecomposable permutation $kG$-module $M$ with stabilizer containing $H$ but strictly contained in $L$. Therefore, $\Psi^{K}(M)=0$, so the image of $\mathbb{F}^{K}(\iota^{H}_{C})$ in degree $h_{C}(H)$ is the zero map, as desired. An analogous argument demonstrates $\mathbb{F}^{H}(\iota^{K}_{C})$ is also the zero map. ∎
Corollary 4.11.
The prime $\mathfrak{m}_{H}$ is the only closed point of $\operatorname{Spc}(\mathscr{K})$ contained in the open $U(H)$.
Proof.
We previously showed in Proposition 4.7 that $\mathfrak{m}_{H}\in U(H)$. Since $U(H)$ is defined by iterating over the irreducible endotrivials, if $h_{C}(H)\neq h_{C}(K)$ for some irreducible endotrivial $C$, then $\mathfrak{m}_{K}\not\in U(H)$. Therefore, it suffices to consider the case where $G$ has a pair of indistinguishable subgroups $H,K$. If this occurs, clearly $H,K<G$, so $H,K<_{G}L\leq G$, where $L$ is the smallest subgroup of $G$ containing both a conjugate of $H$ and $K$. Since $|L|>|H|$, there exists a Borel-Smith function $f$ for which $f(L)\neq f(H)$ (see Remark 4.9). Therefore, there exists a canonical $\mathbb{Z}$-basis element $b$ of $\operatorname{CF}_{b}(G)$ which also satisfies $b(L)\neq b(H)$. Hence, there exists an element $C$ of the canonical $\mathbb{Z}$-basis $\mathcal{B}(G)\cup\{k[1]\}$ of $\operatorname{Pic}(\mathscr{K}(G))$ with $h_{C}(H)\neq h_{C}(L)$. The previous lemma implies $\mathbb{F}^{K}(\iota^{H}_{C})$ is not an isomorphism, and it follows by definition that $\mathfrak{m}_{K}\not\in U(H)$, as desired. ∎
Warning: Corollary 4.11 does not imply that the open $U(H)$ has a unique closed point viewed as a subspace of $\operatorname{Spc}(\mathscr{K})$. The open $U(H)$ has a minimal closed point if and only if $\mathscr{K}(G)|_{U(H)}\cong\mathscr{K}(G)/\mathfrak{m}_{H}$, i.e. $U(H)$ consists of all points specializing to $\mathfrak{m}_{H}$. In general this may not occur.
Remark 4.12.
Corollary 4.11 result does not hold for the open cover of [BG25, Section 13] for non elementary abelian $p$-groups. For instance, [BG25, Proposition 13.14] states the closed complement of the open $U^{\prime}(1)$ (we write $U^{\prime}(1)$ to denote the open $U_{G}(1)$ of [BG25, Proposition 13.14] in order to distinguish the covers apart) is the support of $\operatorname{kos}_{G}(F)$, where $F$ denotes the Frattini subgroup of $G$ (i.e. the intersection of all maximal subgroups of $G$). By [BG25, Corollary 7.17], we have $\mathfrak{m}_{H}\in\operatorname{supp}(\operatorname{kos}_{G}(F))$ if and only if $H\not\leq_{G}F$, hence $\mathfrak{m}_{H}\in U^{\prime}(1)$ if and only if $H\leq_{G}F$. Therefore, $U^{\prime}(1)$ contains a unique closed point of $\mathscr{K}(G)$ if and only if $F=1$ if and only if $G$ is an elementary abelian $p$-group.
Corollary 4.13.
Every element $C\in\operatorname{Pic}(\mathscr{K}(G))$ is a line bundle under the open cover $\{U(H)\}_{H\leq G}$. In particular, for every subgroup $H\leq G$, we have an isomorphism $C\cong k[h_{C}(H)]$ in the localization $\mathscr{K}(G)|_{U(H)}$.
Proof.
By construction of $U(H)$, every irreducible endotrivial corresponding to an element of the canonical $\mathbb{Z}$-basis of $\operatorname{CF}_{b}(G)$ is isomorphic to $k[h_{C}(H)]$ in $\mathscr{K}(G)|_{U(H)}$. As the set of irreducible endotrivials is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, it follows that every endotrivial $C$ is isomorphic to $k[h_{C}(H)].$
Remark 4.14.
Corollary 4.13 does not imply the existence of a homomorphism $f:k[h_{C}(H)]\to C$ which is an isomorphism in $\mathscr{K}(G)|_{U(H)}$. Such homomorphisms may not exist in general for non-effective endotrivials. For example, let $p=2$ and $G=C_{2}$, then the two-term endotrivial $C:=k\xrightarrow{\operatorname{coaug}}kC_{2}$ with $k$ in homological degree 0 satisfies $\operatorname{Hom}_{\mathscr{K}(G)}(k,C)=0$, but we have a local isomorphism $k[0]\cong C$ in the localization $\mathscr{K}(G)|_{U(C_{2})}$.
We turn to our open $U(1).$ First, we need a lemma about faithful endotrivials, i.e. endotrivials that arise from a faithful real representation $V$. Equivalently, these are endotrivials whose h-marks are all 0 for any subgroup $H$ containing a nonzero normal subgroup of $G$, see [Mil24]. In this case, we recover a generalization of [BG25, Proposition 13.14].
Proposition 4.15.
Let $C$ be an effective endotrivial arising from a faithful real representation $V$. Then $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and we have an equality
$\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))=\operatorname{supp}(\ker(\operatorname{Res}^{G}_{1}))=\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{kos}_{G}(1)).$
Proof.
We have that $\ker(V)=1$, so by Proposition 4.1 if $n$ is the highest nonzero homological degree of $C$, $C_{n}\cong kG$. Now consider the complex $\operatorname{cone}(\iota^{1}_{C})^{*}$. After shifting, this complex satisfies $C_{1}\cong kG$ and $C_{0}\cong k$. Therefore, we are in the situation of [BG25, Corollary 3.20], and it follows that $\operatorname{cone}(\iota^{1}_{C})^{*}$, hence $\operatorname{cone}(\iota^{1}_{C})$, generates $\mathscr{K}_{ac}(G)$ as a tt-ideal, and $\operatorname{supp}(\mathscr{K}_{ac}(G))=\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))$, as desired. ∎
Remark 4.16.
Analogously, if $V$ is a real representation with kernel $N\trianglelefteq G$, then it is a faithful real representation of $G/N$, and the associated endotrivial $C$ is a faithful endotrivial complex of $k[G/N]$-modules. It follows that in this case, $\operatorname{cone}(\iota^{1}_{C})=\operatorname{cone}(\iota^{N}_{C})$ (as they are the same map) and $\operatorname{supp}(\ker(\operatorname{Res}^{G}_{N}))=\operatorname{supp}(\operatorname{kos}_{G}(N))$, recovering [BG25, Lemma 13.2].
Theorem 4.17.
Let $G$ be a finite $p$-group. The closed complement of the open $U(1)$ is the support of $\operatorname{kos}_{G}(1)$, i.e. the closed support of the tt-ideal $\mathscr{K}_{ac}(G)=\ker(\operatorname{Res}^{G}_{1})$. In particular, $U(1)$ is equal to the cohomological open $V_{G}=\operatorname{Spc}(\operatorname{D}_{b}(kG)).$
Proof.
Since $\mathcal{B}(G)\cup\{k[1]\}$ is a basis of $\operatorname{Pic}(\mathscr{K}(G))$, for any $\mathscr{P}\in U(1)$, all endotrivials are isomorphic to a shift of the tensor unit in $\mathscr{K}(G)/\mathscr{P}$. In particular, there exists a faithful real representation $V$ (for instance the regular representation $\mathbb{R}G$) and a corresponding endotrivial $C_{V}$, for which $\iota_{C_{V}}^{1}\colon k[h_{C_{V}}(1)]\to C_{V}$ is an isomorphism in $\mathscr{K}(G)/\mathscr{P}$ (in fact, an isomorphism in $\mathscr{K}|_{U(1)}$ by Corollary 4.13). Therefore, $\operatorname{open}(\operatorname{cone}(\iota^{1}_{C_{V}}))\supseteq U(1)$. By Proposition 4.15, we have $\operatorname{supp}(\operatorname{kos}_{G}(1))=\operatorname{supp}(\operatorname{cone}(\iota_{C_{V}}^{1}))\subseteq U(1)^{c}$.
Conversely, by definition we have
$U(1)=\bigcap_{C\in\mathcal{B}(G)}\operatorname{open}(\iota^{1}_{C}).$
Therefore by Remark 4.16,
$U(1)^{c}=\bigcup_{C\in\mathcal{B}(G)}\operatorname{supp}(\operatorname{cone}(\iota^{1}_{C}))=\bigcup_{C\in\mathcal{B}(G)}\operatorname{supp}(\operatorname{kos}_{G}(\ker(C))),$
where $\ker(C)$ denotes the kernel of the real representation corresponding to $C$. For any subgroup $H\leq G$, one has by [BG25, Corollary 7.17] $\operatorname{supp}(\operatorname{kos}_{G}(H))\subseteq\operatorname{supp}(\operatorname{kos}_{G}(1))$, thus $U(1)^{c}\subseteq\operatorname{supp}(\operatorname{kos}_{G}(1))$, thus equality holds. ∎
Corollary 4.18.
We have an tt-equivalence $\mathscr{K}(G)|_{U(1)}\cong\operatorname{D}_{b}(kG)$.
Proof.
This follows immediately from Proposition 4.17 and general tt-geometry. ∎
In particular, the open $U(1)$ has a unique closed point as a subspace of $\mathscr{K}(G)$. However, we stress that in general the other opens $U(H)$ need not have unique closed points.
5. The (re-)twisted cohomology ring
We are now ready to define the twisted cohomology ring for finite $p$-groups. When $G$ is a finite elementary abelian $p$-group, we recover the previously constructed twisted cohomology ring, since in this case the set of endotrivials $\{u_{N}\}_{N\in\mathcal{N}}$ as defined in [BG25, Example 12.1] and the shift of the tensor unit $k[1]$ form precisely the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$.
Definition 5.1.
Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e. tuples of non-negative integers indexed by the $\mathcal{B}(G)$. Equivalently, the monoid of twists identifies with the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$. Consider the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
$\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)=\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G;k):=\bigoplus_{s\in\mathbb{Z}}\bigoplus_{q\in\mathbb{N}^{\mathcal{B}(G)}}\operatorname{Hom}_{\mathscr{K}(G)}\left(k,\bigotimes_{C\in\mathcal{B}(G)}C^{\otimes q(C)}[s]\right).$
Its multiplication is induced by the tensor product in $\mathscr{K}(G)$. Note that only non-positive shifts $s\leq 0$ produce non-zero homomorphisms. We call $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ the twisted cohomology ring of $G$. It is convenient to write
$k(q):=\bigotimes_{C\in\mathcal{B}(G)}C^{q(C)}$
for every twist $q\in\mathbb{N}^{\mathcal{B}(G)}$ and thus abbreviate $\operatorname{H}^{s,q}(G)=\operatorname{Hom}_{\mathscr{K}(G)}(k,k(q)[s])$.
This ring is commutative for $p=2$ and graded-commutative for $p$ odd (see [BG25, Remark 12.8]) - note that only the shift $s$ plays a role in the graded commutativity, and not the twist $k(q)$.
Remark 5.2.
This construction may not immediately adapt if $G$ is not a $p$-group, since in this setting it is not known if one has a canonical $\mathbb{Z}$-basis for $\operatorname{Pic}(\mathscr{K}(G))$. A possible replacement could be the subgroup of $\mathscr{K}(G)$ of endotrivials arising from representation spheres for $G$. We propose that this subgroup has finite index in $\mathscr{K}(G),$ hence the analogous construction of an open cover should also hold.
To examine the $k$-vector space $\operatorname{H}^{s,q}(G)$, it suffices to choose an indecomposable representative due to the stabilizers grow property.
Proposition 5.3.
Let $G$ be a finite $p$-group, $k(q)=C$ an effective endotrivial, and $s$ an integer. Let $f\colon k\to C[s]$ be a nonzero chain complex homomorphism. We have that $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$ if and only if $\operatorname{im}(f)\subset C_{s}$ is contained in a contractible direct summand of $C$. In particular, if $C$ is indecomposable, then
$\operatorname{H}^{s,q}(G):=\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=\operatorname{Hom}_{\operatorname{Ch}(p-\operatorname{perm}(kG))}(k,C[s]).$
Proof.
The converse implication is straightforward. Suppose $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$, then there exists a homotopy $h\colon k\to C_{s+1}$ such that $f=d_{s+1}\circ h$. Choose a canonical $\mathbb{Z}$-basis of $C$, and consider the projection of the image of $f$ onto each indecomposable summand. Necessarily on each summand $k[G/H]$ whose projection is nonzero, the image must be the unique minimal nonzero submodule $k\subseteq k[G/H]$. Since $f=d_{s+1}\circ h$, there exists an indecomposable direct summand $M$ of $C_{s+1}$ such that $p_{k[G/H]}\circ f=d_{s+1}\circ p_{M}\circ h$. Since the stabilizers grow condition holds for $C$, the only possible homomorphisms $M\to k[G/H]$ with $k\subset M$ not in the kernel are isomorphisms, hence $M\cong k[G/H]$ and we may split off the contractible summand. An induction argument demonstrates $\operatorname{im}(f)$ is contained in a contractible direct summand of $C$ as desired. ∎
Just like in [Bal10a, BG25], we have a canonical comparison map.
Proposition 5.4.
There is a continuous comparison map
$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
mapping a tt-prime $\mathscr{P}$ to the ideal generated by the homogeneous $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ whose cone does not belong to $\mathscr{P}$. It is characterized by the fact that for all $f$,
$\operatorname{comp}_{G}^{-1}(Z(f))=\operatorname{supp}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is not invertible in }\mathscr{K}(G)/\mathscr{P}\}$
where $Z(f)=\{\mathfrak{p}\mid f\in\mathfrak{p}\}$ is the closed subset of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ defined by $f$.
Proof.
Same as [BG25, Proposition 13.4]. ∎
Recall the conservative functor of Theorem 1.8, the collection of modular fixed points functors $\{\hat{\Psi}^{H}\}_{H\leq G}$. For shorthand, we denote this functor by $\prod\hat{\Psi}:\mathscr{K}(G)\to\prod_{H\leq G}\operatorname{D}_{b}({G}/\!\!/{H})$. We’ll describe the “twisted” analogue of the induced homomorphism on the respective cohomology rings.
Definition 5.5.
Let $C\in\mathcal{B}(G)$, $s\in\mathbb{Z}$, and $f\colon k\to C[s]$ be a homogeneous element of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. For every subgroup $H\leq G$, $\hat{\Psi}^{H}(C)\in\operatorname{Hom}_{\operatorname{D}_{b}({G}/\!\!/{H})}(k,\Psi^{H}(C)[s])$, and since we have an isomorphism $\Psi^{H}(C)[s]\cong k[h_{C}(H)+s]$ (determined by $\iota^{H}_{C}$), this determines an element in group cohomology $f_{H}\in\operatorname{H}^{h_{C}(H)+s}({G}/\!\!/{H},k)=\operatorname{Ext}^{h_{C}(H)+s}_{k[{G}/\!\!/{H}]}(k,k)$ determined by the fraction
$k\xrightarrow{\Psi^{H}(f)}C[s]\xleftarrow{\Psi^{H}(\iota^{H}_{C})}k[h_{C}(H)+s].$
We consider $\hat{\Psi}^{H}(f)$ an element of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$, i.e. as a morphism in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$. Note that for the endotrivial $u_{N}$ defined in [BG25, Definition 12.3] and $H=1$, the fraction $\hat{\Psi}^{1}(\iota^{G}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{+}$ of [BG25, Remark 12.7]. If $H=G$, the fraction $\hat{\Psi}^{G}(\iota^{1}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{-}$.
Performing this construction over all subgroups $H\leq G$ up to conjugacy determines an element $\left(\prod\hat{\Psi}\right)(f)$ in the $\operatorname{CF}(G)$-graded ring
$\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G):=\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k).$
For $\prod\hat{\Psi}$ to be a well-defined ring homomorphism, the denominator associated to an arbitrary effective endotrivial $C$ is defined as the product of $\iota^{H}_{C}$s corresponding to the unique tensor product factorization of $C$ into a product of irreducible endotrivials. This product of $\iota^{H}$s remains a quasi-isomorphism. Note that this product is well-defined up to reordering, since $h_{C}$ returns only even values for $p$ odd.
The assignment described describes the ring homomorphism $\prod\hat{\Psi}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{dH}^{\scriptscriptstyle\bullet}(G)$ induced by the conservative functor of [BG25, Theorem 7.2]. It is clear that this homomorphism is linear with respect to addition. Linearity over multiplication follows from the following roof; let $C$ and $D$ be endotrivials, $H$ a subgroup of $G$, $i$ and $j$ integers, and $f\colon k\to C[i]$ and $g\colon k\to D[j]$ homogeneous elements of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. For shorthand, we write $\iota^{H}_{C}$ and $\iota^{H}_{D}$ for the products of $\iota^{H}_{C^{\prime}}$ and $\iota^{H}_{D^{\prime}}$ running over all irreducible endotrivial $C^{\prime}$ and $D^{\prime}$ occurring in the tensor decompositions of $C$ and $D$ respectively. Then, the following roof commutes.
[…]
Note that $\prod\hat{\Psi}$ sends tensor products of homomorphisms to Yoneda products of cohomology classes, sums of homomorphisms to Baer sums, and sends a homogeneous element in homological degree $(s,k(q))$ to a homogeneous element in degree $\underline{s}+h_{k(q)}$, where $\underline{s}$ denotes the constant superclass function returning only $s$.
Finally, we of course have a homeomorphism
$\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}V_{{G}/\!\!/{H}}.$
We denote the induced map on spectrum by $\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$, similar to the categorical case.
Proposition 5.6.
The kernel of $\prod\hat{\Psi}$ is $\otimes$-nilpotent, i.e. elements $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ for which $f^{\otimes n}=0$ for some $n\geq 1$.
Proof.
This follows directly from [BG25, Theorems 7.1, 7.2], since $\prod\hat{\Psi}$ is induced from the conservative family of functors $\{\hat{\Psi}^{H}\}$. ∎
Proposition 5.7.
The following diagram in $\operatorname{Top}$ commutes.
[…]
In particular, the image of the comparison map $\operatorname{comp}_{G}$ is as a set, precisely the image of $\prod\hat{\psi}$. Moreover, the image of $\prod\hat{\psi}$ is dense in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$.
Proof.
Every prime $\mathscr{P}(H,\mathfrak{p})\in\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ (in the notation of [BG25]) is explicitly the product of the prime $\mathfrak{p}$ of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ and the full rings $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{K})$ with $K$ different from $H$. The map $\prod\hat{\psi}$ sends the prime ideal $\mathscr{P}(H,\mathfrak{p})\in V_{{G}/\!\!/{H}}\subseteq\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))$ to the ideal $\{f\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))\mid\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})\}$, and we have $\left(\prod\hat{\psi}\right)(f)\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$. Verification of commutativity now follows from the fact that given any $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$, $H\leq G$, and $\mathfrak{p}\in V_{{G}/\!\!/{H}}$, we have $\operatorname{cone}(f)\not\in\mathscr{P}(H,\mathfrak{p})$ if and only if $\hat{\Psi}^{H}(f)\in\mathfrak{p}$. Indeed, it is well-known that the comparison map on the tt-category $\operatorname{D}_{b}({G}/\!\!/{H})$ is a homeomorphism $\operatorname{Spc}(\operatorname{D}_{b}({G}/\!\!/{H}))\xrightarrow{\cong}\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$, sending a prime $\mathscr{P}(H,\mathfrak{p})$ to the prime $\mathfrak{p}=\{\zeta\mid\operatorname{cone}(\zeta)\not\in\mathfrak{p}\}\subset\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k)$. Thus, $\zeta\in\mathfrak{p}$ if and only if $\operatorname{cone}(\zeta)\not\in\mathscr{P}(H,\mathfrak{p})$, and setting $\zeta=\hat{\Psi}^{H}(f)$ shows the result.
For the final statement, Proposition 5.6 implies the kernel of $d$ is nilpotent, so the result follows by a standard algebraic geometry result. ∎
Remark 5.8.
In particular, the image of the prime $\mathscr{P}(H,\mathfrak{p})$ under the comparison map $\operatorname{comp}_{G}$ is equivalently
$\left(\prod\hat{\psi}\right)(\mathscr{P}(H,\mathfrak{p}))=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\hat{\Psi}^{H}(f)\in\mathfrak{p}\}.$
5.1. Twisted cohomology under localization and tt-functors
Next, we’ll extend the results of [BG25, Section 14] regarding localization.
Definition 5.9.
Let $H$ be a subgroup of $G$. Let $S_{H}\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ be the multiplicative subset generated by all $\iota^{H}_{C}$, where $C$ runs over the irreducible endotrivials $C\in\mathcal{B}(G)$. We define a $\mathbb{Z}$-graded ring
$\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H):=\big{(}\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)[S_{H}^{-1}]\big{)}_{\text{0-twist}}$
as the twist-zero part of the localization of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with respect to $S_{H}$. Explicitly, the homogeneous elements of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ consist of fractions $\frac{f}{g}$ where $f,g\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with the same $\mathcal{B}(G)$-twist $q$, and $g$ is a product of “$H$-local isomorphisms” $g^{\prime}\colon k\to C[s]$ with $C\in\mathcal{B}(G)$. Thus, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is $\mathbb{Z}$-graded by the shift only. The homological degree of $\frac{f}{g}$ is the difference sts-t between the shifts of $f$ and $g$. In particular, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(1)\cong\operatorname{H}^{\scriptscriptstyle\bullet}(G)$, following directly from Definition 5.5.
Remark 5.10.
Recall for a morphism $f$ in $\mathscr{K}(G)$ we write
$\operatorname{open}(f):=\operatorname{open}(\operatorname{cone}(f))=\{\mathscr{P}\mid f\text{ is invertible in }\mathscr{K}(G)/\mathscr{P}\}.$
We have that $\operatorname{open}(f)$ is the preimage of $\operatorname{comp}_{G}$ of the principle open $Z(f)^{c}=\{\mathfrak{p}\mid f\not\in\mathfrak{p}\}$, and is the open locus of $\operatorname{Spc}(\mathscr{K}(G))$ where $f$ is invertible.
Construction 5.11.
We can perform the central localization (see [Bal10a]) of the whole category $\mathscr{K}(G)$
$\mathcal{L}_{G}(H):=\mathscr{K}(G)[S_{H}^{-1}].$
In fact, this localization has idempotent completion $\mathscr{K}(G)|_{U(H)}$ by the same argument as in [BG25, Construction 14.12] or [Bal10a, Theorem 3.6]. Explicitly, the category $\mathcal{L}_{G}(H)$ is the Verdier quotient of $\mathscr{K}(G)$ by the tt-ideal $\langle\{\operatorname{cone}(g)\mid g\in S_{H}\}\rangle$. It has the same objects as $\mathscr{K}(G)$ and morphisms $x\to y$ of the form $\frac{f}{g}$ where $g\colon k\to C$ belongs to $S_{H}$ for $C$ an effective endotrivial and $f\colon x\to C\otimes y$ is any morphism in $\mathscr{K}(G)$ with the same twist $C$ as the denominator $g$. Moreover, the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathcal{L}_{G}(H)}(k)$ of the unit in $\mathcal{L}_{G}(H)$ is the $\mathbb{Z}$-graded ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$.
Construction 5.12.
Twisted cohomology $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is graded over a monoid over the form $\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}$. The ring homomorphisms induced by tt-functors or localization will be homogeneous with respect to a certain homomorphism $\gamma$ on the corresponding grading monoids.
Let $H\leq G$ be a subgroup and consider the central localization $(-)_{U(H)}\colon\mathscr{K}(G)\twoheadrightarrow\mathcal{L}_{G}(H)$. Here, the morphisms $\iota^{H}_{C}$ become isomorphisms, yielding a homomorphism on the grading
$\gamma=\gamma_{U(H)}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}$
defined by $\gamma(s,q)=s+h_{k(q)}(H)$ and we obtain a ring homomorphism
$(-)_{U(H)}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{End}^{{\scriptscriptstyle\bullet}}_{\mathcal{L}_{G}(H)}(k)=\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H),$
homogeneous with respect to the homomorphism $\gamma_{U(H)}$.
We may also consider the effects of restriction and modular fixed points on twisted cohomology. Note $\mathcal{B}(G)$ is more complicated than $\mathcal{N}\subseteq\mathcal{B}(G)$, where $\mathcal{N}$ consists of endotrivials inflated from normal subgroups of $G$ of index $p$, so in general, the shifts $\gamma$ may be more difficult to compute in the case of restriction.
Construction 5.13.
Let $H\trianglelefteq G$ be a normal subgroup. The tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}(G/H)$ maps effective endotrivials with h-mark at $G$ equal to 0 to effective endotrivials at $G/H$ equal to 0. In fact, since deflation preserves irreducibility of a representation, $\Psi^{H}$ sends elements of $\mathcal{B}(G)$ to $\mathcal{B}(G/H)$.
This defines a homomorphism of graded monoids
$\gamma=\gamma_{\Psi^{H}}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G/H)}$
given by $\gamma(s,q)=(s,\overline{q})$ where $\overline{q}$ is given by the surjection $\mathbb{N}^{\mathcal{B}(G)}\twoheadrightarrow\mathbb{N}^{\mathcal{B}(G/H)}$ along the inclusion $\mathcal{B}(G/H)\hookrightarrow\mathcal{B}(G)$ induced from inflation. Therefore, modular fixed points defines a ring homomorphism $\Psi^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ homogeneous with respect to $\gamma_{\Psi^{H}}$.
Given a group homomorphism $\alpha:G^{\prime}\to G$, restriction along $\alpha$ also defines a tt-functor $\alpha^{*}:\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$. This again defines a corresponding ring homomorphism $\alpha^{*}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ homogeneous with respect to $\gamma_{\alpha^{*}}$.
Note in this case if $\alpha$ is not surjective, then $\gamma_{\alpha^{*}}$ may not be constant on the shift $s$ and in general $\overline{q}$ will be difficult to compute. Conversely, if $\alpha$ is surjective, then $\alpha^{*}$ is simply inflation, which may be easily computed, as it is section of modular fixed points. It follows that the homomorphism $\Psi^{H}$ on twisted cohomology is split surjective.
Remark 5.14.
Suppose $F\colon\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$ is a tt-functor and the induced homomorphism $F\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ is homogeneous with respect to $\gamma=\gamma_{F}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G^{\prime})}$ (for instance, modular fixed points or restriction). Then the following square commutes, since $F(\operatorname{cone}(f))=\operatorname{cone}(F(f))$.
$\operatorname{Spc}(\mathscr{K}(G^{\prime}))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{F^{*}}$$\scriptstyle{\operatorname{comp}_{G^{\prime}}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{F^{*}}$
In particular, setting the functor $F$ equal to the localization functor $(-)|_{U(H)}$ obtains the following commutative square.
$\operatorname{Spc}(\mathcal{L}_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\mathcal{L}_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$
Here, the left hand vertical map is the usual twist-free comparison map from [Bal10a] for the tt-category $\mathcal{L}_{G}(H)$ and the $\otimes$-invertible $k[1]$.
We combine the functors as before to strengthen [BG25, Proposition 14.21].
Proposition 5.15.
Let $H\trianglelefteq G$ be a normal subgroup. Then we have a commutative square
$V_{G/H}=\operatorname{Spc}(\operatorname{D}_{b}(k[G/H]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}(G/H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[G/H])}}$$\scriptstyle{\operatorname{comp}_{G}}$
In particular, the diagonal is injective.
Proof.
The proof follows essentially the same as the proof of [BG25, Proposition 14.21], except without the added assumption that $G/H$ is elementary abelian. For this, we use Proposition 4.17 which asserts the quotient $\mathscr{K}(G/H)\to\operatorname{D}_{b}(k[G/H])$ is the central localization $(-)|_{U(1)}$ for any group $G/H$. ∎
6. Injectivity of the comparison map
The crux of injectivity of the comparison map is showing that, the diagonal of Proposition 5.15 is injective when $H$ is not necessarily normal in $G$ as well.
Theorem 6.1.
Let $G$ be a finite $p$-group. The comparison map
$\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
of Proposition 5.4 is injective.
Proof.
Let $\mathscr{P}=\mathscr{P}(H_{1},\mathfrak{p})$ and $\mathscr{Q}=\mathscr{P}(H_{2},\mathfrak{q})$ in $\operatorname{Spc}(\mathscr{K}(G))$ satisfy $\operatorname{comp}_{G}(\mathscr{P})=\operatorname{comp}_{G}(\mathscr{Q})$ in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. This implies that $\mathscr{P}\in\operatorname{open}(f)$ if and only if $Q\in\operatorname{open}(f)$ for every $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. In particular, for every effective endotrivial $C\in\mathcal{B}(G)$ and subgroup $H\leq G$, we have that $\mathscr{P}\in\operatorname{open}(\iota^{H}_{C})$ if and only if $\mathscr{Q}\in\operatorname{open}(\iota^{H}_{C})$.
Suppose for contradiction that $H_{2}\not\leq_{G}H_{1}$. By Corollary 4.11, there exists an effective endotrivial $C\in\mathcal{B}(G)$ for which $\Psi^{H_{2}}(\iota_{C}^{H_{1}})$ is the zero map, and vice versa. Now, consider the map $\hat{\psi}^{H_{2}}\colon V_{{G}/\!\!/{H_{2}}}\hookrightarrow\operatorname{Spc}(\mathscr{K}(G))$. We have:
$(\hat{\psi}^{H_{2}})^{-1}(\operatorname{open}(\iota^{H_{1}}_{C}))$ $=\operatorname{open}(\operatorname{cone}(\hat{\Psi}^{H_{2}}(\iota^{H_{1}}_{C})))$
$=\operatorname{open}(0\colon k\to\hat{\Psi}^{H_{2}}(C))$
$=\emptyset$
Therefore, $V_{{G}/\!\!/{H_{2}}}\cap\operatorname{open}(\iota^{H_{1}}_{C})=\emptyset$ in $\operatorname{Spc}(\mathscr{K}(G))$. On the other hand, $V_{{G}/\!\!/{H_{1}}}\subseteq\operatorname{open}(\iota^{H_{1}}_{C})$, since $\iota^{H_{1}}_{C}$ is invertible in $\mathscr{K}(G)/\mathfrak{m}_{H_{1}}$. Therefore, $\mathscr{P}\in\operatorname{open}(\iota^{H_{1}}_{C})$ but $\mathscr{Q}\not\in\operatorname{open}(\iota^{H_{1}}_{C})$, a contradiction. Thus $H_{2}\leq_{G}H_{1}$, and by symmetry $H_{1}=_{G}H_{2}$. Set $H:=H_{1}.$
If $H\trianglelefteq G$, then we have two points $\mathfrak{p},\mathfrak{q}\in V_{{G}/\!\!/{H}}$ that go to the same image under
$V_{{G}/\!\!/{H}}\xrightarrow{\hat{\psi}^{H}}\operatorname{Spc}(\mathscr{K}(G))\xrightarrow{\operatorname{comp}_{G}}\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
and this map is injective from Proposition 5.15. Otherwise, we have the following commutative diagram.
$V_{N_{G}(H)/H}$$\operatorname{Spc}(\mathscr{K}(N_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{Res}^{*}}$$\scriptstyle{\operatorname{comp}_{N_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\operatorname{Res}^{*}}$
Here, the diagonal arrow is injective since $H\trianglelefteq N_{G}(H)$, and the top row is injective since $\operatorname{Spc}(\mathscr{K}(G))=\bigsqcup_{H\in s_{p}(G)/G}V_{{G}/\!\!/{H}}$, and the composition of the above maps is precisely $\hat{\psi}^{H}$. Therefore, to show injectivity, it suffices to show the bottom composition of maps
$V_{N_{G}(H)/H}\hookrightarrow\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective. We establish this as a separate lemma.
Lemma 6.2.
Let $H$ be a subgroup of $G$. The composition of continuous maps
$V_{N_{G}(H)/H}\hookrightarrow\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(N_{G}(H)))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective.
Proof.
Since $G$ is a finite $p$-group, we may choose a subnormal sequence $N_{G}(H)=H_{0}\triangleleft H_{1}\triangleleft\cdots\triangleleft H_{n}=G$ with $H_{i+1}/H_{i}=C_{p}$. We have the following setup depicted by the commutative diagram below.
$V_{N_{G}(H)/H}$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{0}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{1}))$$\cdots$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{n}))$}
Here, the horizontal arrows are induced by restriction and the downwards arrows are $\hat{\psi}^{H}$. We will prove this statement inductively by assuming that the map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ is injective and showing that the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, which then of course implies the next map $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective as well. Injectivity of the base case $H_{0}=N_{G}(H)$ is established by Proposition 5.15.
Let $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$. We claim that $\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{p})=\operatorname{Spec}(\operatorname{Res}^{H_{i+1}}_{H_{i}})(\mathfrak{q})$ if and only if the following property $(\ast)$ is satisfied:
$(\ast)$  For every $f\in\mathfrak{p}$, there exists a $g\in H_{i+1}$ such that $^{g}f\in\mathfrak{q}$.
The converse is straightforward. Suppose there exists an $f\colon k\to C[s]\in\mathfrak{p}$ for which $^{g}f\not\in\mathfrak{q}$ for all $g\in H_{i+1}$. Then, the trace product of $H_{i+1}$-conjugates
$f^{\prime}:=\prod_{g\in H_{i+1}/H_{i+1}}{}^{g}f$
is a morphism $f^{\prime}\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1})$. Indeed, the tensor product of chain complexes
[…]
has $kH_{i+1}$-module structure (one may see for instance that $C^{\prime}$ is the restriction of the tensor induced complex $\operatorname{Ten}^{H_{i+1}}_{H_{i}}C$ via the Mackey formula - while tensor induction is not defined in general up to homotopy, in this case the restriction is well-behaved up to homotopy), and the image of $k$ in $C^{\prime}$ is $H_{i+1}$-stable. By primality of $\mathfrak{p}$ and $\mathfrak{q}$, we have $f^{\prime}\in\mathfrak{p}$ but $f^{\prime}\not\in\mathfrak{q}$, as desired.
Now, we show that if $\mathfrak{p},\mathfrak{q}\in\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))$ satisfy $(\ast)$ and belong to the image of $V_{N_{G}(H)/H}$, then $\mathfrak{p}=\mathfrak{q}$. The statement is trivial if $H_{i+1}/H_{i}$ acts on $H_{i}$ trivially, so assume the action is nontrivial. Recall from Proposition 5.7 that in this case, $\mathfrak{p}=\rho(H,\overline{\mathfrak{p}})$ with $\overline{\mathfrak{p}}\in V_{N_{G}(H)/H}$, and we have $\mathfrak{p}=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i})\mid\hat{\Psi}^{H}(f)\in\overline{\mathfrak{p}}\}$. Let $f\in\mathfrak{p}$. By Lemma 4.10, there exists an effective endotrivial chain complex $C$ of $kH_{i}$-modules and morphism $\iota^{H}_{C}$ such that $\hat{\Psi}^{H}(\iota^{H}_{C})=\operatorname{id}_{k}$ and $\hat{\Psi}^{H}({}^{g}(\iota^{H}_{C}))=0$ for all $g\in H_{i+1}\setminus H_{i}$. Therefore, $\iota^{H}_{C}\not\in\mathfrak{p}$ (since $\overline{\mathfrak{p}}\neq V_{N_{G}(H)/H}$), but for all $g\in H_{i+1}\setminus H_{i}$, we have $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{p}$. We construct a morphism $f^{\prime\prime}\in\mathfrak{p}$ as follows,
$f^{\prime\prime}:=f+\left(\sum_{g\in[(H_{i+1}/H_{i})-H_{i}]}{}^{g}(\iota^{H}_{C})\cdot(\operatorname{id}_{k}-f)\right)\in\mathfrak{p}.$
By construction $\hat{\Psi}^{({}^{g}H)}(f^{\prime\prime})$ is an isomorphism for all $g\in H_{i+1}\setminus H_{i}$, therefore $^{g}(f^{\prime\prime})\not\in\mathfrak{q}$. But since $\mathfrak{p},\mathfrak{q}$ satisfy $(\ast)$, it follows that $f^{\prime\prime}\in\mathfrak{q}$ as well. Since each $^{g}(\iota^{H}_{C})\in\varphi(H,0)\subseteq\mathfrak{q}$, cancellation implies $f\in\mathfrak{q}$ as well. Thus $\mathfrak{p}\subseteq\mathfrak{q}$, and symmetry implies $\mathfrak{p}=\mathfrak{q}$, and the composition $V_{N_{G}(H)}\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i}))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H_{i+1}))$ is injective, as desired. ∎
As an immediate corollary, we obtain a strengthening of Proposition 5.15.
Corollary 6.3.
Let $H$ be any subgroup of $G$. Then we have a commutative square
$V_{{G}/\!\!/{H}}=\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\hat{\psi^{H}}}$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}}$$\scriptstyle{\operatorname{comp}_{G}}$
In particular, the diagonal is injective, and the product
$\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$
is injective.
Corollary 6.4.
If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, then for any $H\leq G$, the comparison map restricts to a homeomorphism
$\operatorname{comp}_{G}\colon U(H)\cong\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)).$
In particular, $\operatorname{comp}_{G}\colon\operatorname{Spc}(\mathscr{K}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is an open immersion, with image the open subspace of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ with closed points (maximal primes)
$\operatorname{comp}_{G}(\mathfrak{m}_{H})=\{f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\mid\Psi^{H}(f)\text{ is not a quasi-isomorphism}\}.$
Proof.
We have already an injective map by Remark 5.14 and Theorem 6.1
$\operatorname{comp}_{\mathscr{L}_{G}(H)}:\operatorname{Spc}(\mathscr{L}_{G}(H))=U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))).$
Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ implies $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is noetherian as well, hence [Bal10a, Theorem 7.3] implies $\operatorname{comp}_{\mathscr{L}_{G}(H)}$ is a continuous bijection. It remains to show this map is a homeomorphism; we show it is closed. Let $\mathscr{L}_{G}(H)_{\langle k\rangle}$ denote the unitization of $\mathscr{L}_{G}(H)$, i.e. the tt-subcategory compactly generated by the tensor unit. Of course, the unitization has the same cohomology ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$. Then [San25, Corollary 10.2] asserts that the inclusion $\mathscr{L}_{G}(H)_{\langle k\rangle}\to\mathscr{L}_{G}(H)$ induces a homeomorphism $\operatorname{Spc}(\mathscr{L}_{G}(H)_{\langle k\rangle})\cong\operatorname{Spc}(\mathscr{L}_{G}(H))$. The comparison map factors through $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}$ as follows:
$\operatorname{Spc}(\mathscr{L}_{G}(H))\to\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)).$
Now, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is (by definition) generated by its tensor unit. By noetherianity of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$, $\mathscr{L}_{G}(H)_{\langle k\rangle}$ is therefore End-finite in the sense of [Lau23, Definition 2.6], hence [Lau23, Proposition 2.7] asserts the induced map on spectra $\operatorname{Spc}(\mathscr{L}_{G}(H))_{\langle k\rangle}\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism, thus $\operatorname{comp}_{\mathscr{L}_{G}(H)}:U(H)\to\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$ is a homeomorphism as well, as desired.
Therefore, $\operatorname{comp}_{G}$ is a homeomorphism onto its image in $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. It is easy to check the closed points $\mathfrak{m}_{H}$ are mapped to the corresponding maximal primes in the image as described in the theorem statement. ∎
The following corollary is immediate.
Corollary 6.5.
Let $\mathcal{O}_{G}^{\scriptscriptstyle\bullet}$ denote the sheaf of $\mathbb{Z}$-graded rings on $\operatorname{Spc}(\mathscr{K}(G))$ obtained by sheafifying $U\mapsto\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathscr{K}(G)|_{U}}(k)$. If $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian, $(\operatorname{Spc}(\mathscr{K}(G)),\mathcal{O}_{G}^{\scriptscriptstyle\bullet})$ is a Dirac scheme.
Remark 6.6.
We informally say some words about noetherianity to close, to hopefully conjure hope within the reader that for any finite $p$-group $G$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian. First, one can show $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian when $G$ satisfies the following property: the indecomposable representative of each irreducible endotrivial $C$ in $\mathcal{B}(G)$ has at most one indecomposable module in every homological degree. This holds for instance (and possibly only when) $G$ is abelian, where the endotrivials correspond to subgroups $N\leq G$ such that $G/N$ is cyclic (see [Mil24, Section 6]). With this property, the proof follows essentially identically as [BG25, Lemma 12.12], with that the maps $\iota^{1}_{C}$ and $\iota^{G}_{C}$ replacing the maps $b_{N}$ and $a_{N}$ respectively for each irreducible endotrivial $C\in\mathcal{B}(G)$.
One crux of the matter is that if an irreducible endotrivial has a homological degree with two or more indecomposables, the exact sequence [BG25, 12.14] sends a term $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,v\otimes u^{\otimes q})$ to a direct sum of morphisms $(f_{1},\dots,f_{n})$ with $f_{i}\in\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$, where the subgroups $H_{i}\leq G$ are the corresponding stabilizers of the indecomposable permutation modules. It is unclear if one can inductively lift this sum of morphisms to a morphism in $\operatorname{Hom}_{\mathscr{K}(G)}(1,v\otimes u^{\otimes q})$, polynomial in morphisms $k\to C[s]$ with $C\in\mathcal{B}(G)$, such that its image in every $\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$ corresponds to the image of $f$. Additionally, the lengths of the irreducible endotrivials $C$ can be arbitrarily long, but an inductive argument may be used to resolve this issue.
A potential roadmap to proving noetherianity modulo nilpotents is to verify that the image of the map $\prod\hat{\Psi}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ is finite, as the kernel of $\prod\hat{\Psi}$ is nilpotent (Proposition 5.6). One can further reduce this to the elementary abelian case by postcomposing by restriction to all elementary abelian subquotients of $G$, as the restriction $\operatorname{H}^{\scriptscriptstyle\bullet}(G)\to\prod_{E\in\operatorname{elemab}}(G)\operatorname{H}^{\scriptscriptstyle\bullet}(E)$ detects nilpotents as well (see [Ben98b, Proposition 5.2.2]). It is not too unreasonable to expect this to hold - for instance [Eve61, Theorem 7.1] states that for any subgroup $H\leq G$, $\operatorname{H}^{\scriptscriptstyle\bullet}(H)$ is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$ via restriction. Hence $\operatorname{H}^{\scriptscriptstyle\bullet}(H)^{G}$ (i.e. the image of restriction, see [Ben98a, Proposition 3.8.2]), is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$. For this reason, it is quite pertinent and interesting to determine the image of $\hat{\Psi}^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ for any subgroup $H\leq G$. We note for instance that one can show the image of $\prod\hat{\Psi}$ is finite when $p=2$ and $G=Q_{8}$ exploiting the fact that every subgroup of $Q_{8}$ is normal. However, it is not clear if an analogue of [Bal10a, Theorem 7.3] under the hypothesis that the cohomology ring is noetherian modulo nilpotents can be utilized.